### v1

# The Yōkai Learning Environment: Tracking Beliefs Over Space and Time

Abstract
Developing collaborative AI hinges on Theory of Mind (ToM) – the ability to reason about the beliefs of others to build and maintain common ground. Existing ToM benchmarks, however, are restricted to passive observer settings or lack an assessment of how agents establish and maintain common ground over time. To address these gaps, we introduce the Yōkai Learning Environment (YLE) – a multi-agent reinforcement learning (RL) environment based on the cooperative card game Yōkai. In the YLE, agents take turns peeking at hidden cards and moving them to form clusters based on colour. Success requires tracking evolving beliefs, remembering past observations, using hints as grounded communication, and maintaining common ground with teammates. Our evaluation yields two key findings: First, current RL agents struggle to solve the YLE, even when given access to perfect memory. Second, while belief modelling improves performance, agents are still unable to effectively generalise to unseen partners or form accurate beliefs over longer games, exposing a reliance on brittle conventions rather than robust belief tracking. We use the YLE to investigate research questions in belief modelling, memory, partner generalisation, and scaling to higher-order ToM.
1 Introduction
Effective collaboration among partners requires common ground – the shared knowledge, beliefs, and assumptions that enable them to interpret each other’s actions and respond appropriately Clark (1996). When common ground cannot be reliably maintained, collaboration often breaks down Klein et al. (2005). To establish common ground, humans use Theory of Mind (ToM) Premack and Woodruff (1978) to infer others’ mental states. Consequently, enabling computational agents to perform ToM reasoning has been identified as a critical step toward effective human-AI collaboration Klien et al. (2004); Dafoe et al. (2020, 2021).
Research in cognitive science suggests that human ToM integrates multiple reasoning processes – including memory, spatial navigation, and mentalising Suddendorf and Corballis (2007); Mitchell (2009). Prior work in computational ToM has focused on the latter, for example, by predicting others’ behaviour from an omniscient observer’s perspective Baker et al. (2011); Rabinowitz et al. (2018); Gandhi et al. (2021). However, humans are rarely passive observers of a scene where they solely predict behaviour Sclar et al. (2022). Instead, they are active participants and must adjust their ToM reasoning to the interactive setting in which their own actions change the environment, influence the behaviour of others, and vice versa. This interactive structure necessitates tracking how common ground – the mutually recognised shared experience built through interaction Tomasello (2010) – evolves during collaboration.
The goal of this paper is thus to
propose an environment in which effective collaboration between computational agents requires building common ground by tracking beliefs over space and time.
Figure 1: The YLE poses a challenging ToM reasoning task for cooperative agents in a dynamic environment. Agents cannot observe all cards in a single game. Successful play requires agents to remember and reason about others and their own mental states. In the example, it’s Refer to caption ’s turn. Refer to caption observes the colour of two cards (1 and 2) privately. Refer to caption recalls from earlier that card 1 is blue. When Refer to caption moves card 1 next to card 2, Refer to caption can infer that the second card is likely blue as well. Refer to caption and Refer to caption can now assume this is common ground going forward.
This is the Yōkai Learning Environment (YLE) [^1] Code available at https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env. – a novel multi-agent reinforcement learning (RL) environment based on the collaborative card game Yōkai. In the YLE, players must collaborate to group face-down cards by colour over multiple rounds without direct communication (see Figure 1, left). Each round, one player privately observes the colour of two cards, moves one of them to a new position on the table, and may place a hint card to communicate information about unobserved card colours to others (see Figure 2). Importantly, players may end the game early for a higher reward, but at the risk of losing the game. This makes it a high-stakes, all-or-nothing decision that requires strong ToM reasoning to infer card colours from partners’ actions rather than first-hand observation.
The design of YLE creates a uniquely challenging coordination task. Agents must: (1) remember the colours of moving cards to avoid redundant checks, (2) infer the beliefs and knowledge of others to establish and maintain common ground, (3) interpret teammates’ actions and hints as implicit belief updates, (4) use hint cards as grounded communication, and (5) assess when to end early based on inferred shared knowledge. These challenges require belief tracking over both space (shifting card locations) and time (multi-round gameplay with limited observability), and together reflect key components of ToM Premack and Woodruff (1978); Clark (1996); Pillow (1989); Southgate et al. (2007); Tomasello (2010) in a dynamic, interactive setting. Effectively establishing common ground requires aligning on ad hoc conventions for grouping cards with partners. Prior environments, such as Hanabi Bard et al. (2020), were used to study zero-shot coordination and ToM, but lack the spatial-temporal structure, evolving belief dynamics, and the challenge of ending early based on inferred knowledge.
We evaluate a range of agents on YLE and highlight key limitations in their ability to track common ground in this novel spatial-temporal belief reasoning task. Our experiments show that even with access to perfect memory, agents struggle to collaborate effectively. Among agents without perfect memory, those with explicit memory modules perform best, and performance drops further in four-player settings. While belief representation learning improves performance, agents fail to generalise to new partners or maintain accurate beliefs over longer time horizons. Instead, they rely on brittle conventions rather than robust and generalisable belief tracking. Our contributions are:
  1. 1.
    A fast JAX-based YLE – a new cooperative multi-agent RL benchmark for belief tracking, common ground, and ToM, training at hundreds of thousands of steps per second.
  2. 2.
    An empirical analysis showing that agents fail to coordinate effectively in YLE, and in particular fail to end games early – a behaviour humans achieve reliably.
  3. 3.
    A demonstration that belief representation learning improves performance, but agents still fail to generalise to novel partners or maintain accurate beliefs over time.
  4. 4.
    An evaluation of YLE in a four-player setting, revealing the increased difficulty of maintaining common ground when high-order ToM is required.
2 Yōkai
Figure 2: A sample round in nine-card YLE. Agent Refer to caption first observes two cards (one blue, one green), moves the blue card, and finishes by either (4a) revealing or (4b) placing a hint. Refer to caption places a hint. Then, Refer to caption can start its turn. Refer to caption chooses to end the game. Ending the game will return the final reward based on the outcome. This game is lost because while blue cards are clustered, green and red ones are not.
Yōkai [^2] Detailed rules for Yōkai are available at http://boardgame.bg/yokai%20rules.pdf. is a turn-based cooperative game in which players must sort face-down cards into same-colour clusters to win. A key requirement is that all cards must be connected to at least one other card on their sides. Classically, the game is played by two to four players with 16 cards of four colours each. Complementing these 16 Yōkai cards, additional hint cards feature one or multiple colours and are used to hint at the colour of a face-down card underneath them. They are the only allowed communication mechanism in the game. The number of players determines the number of hint cards, which varies between seven and 10 (see Appendix B for further details). Hint cards are initially face-down and must be revealed before being placed face-up on a Yōkai card. Importantly, this Yōkai card is locked in place and can no longer be moved or observed. Since most hints can be multi-coloured, their interpretation is left to the other agents, depending on the context and history.
A typical turn in Yōkai consists of four steps (see Figure 2). The current player first observes two cards privately, then moves one card, and then either (a) reveals or (b) places a face-up hint card. Consequently, the available legal moves depend on the specific card to be moved. Figure 11, while Figure 10 show legal moves and card configurations. The game ends when all hint cards have been revealed and placed. Optionally, players can end the game prematurely instead of observing cards at the beginning of their turn. The final game score is based on the final card configuration:
$\begin{split}S=&\;5N_{\text{hints\_face\_down}}+2N_{\text{hints\_not\_placed}}+N_{\text{hints\_correct}}\\ &-N_{\text{hints\_wrong}}.\end{split}$ (1)
Players are encouraged to end the game early to achieve a higher score, since unused hint cards contribute most to the reward. However, the earlier a player ends a game, the less information they can acquire first-hand and thus must infer knowledge from others’ behaviour (see Figure 12 for an example). As such, the final game score measures players’ joint ToM reasoning abilities. In particular, the rate at which players successfully end the game early reflects how much they rely on belief inference, making it a powerful new implicit measure of ToM reasoning under uncertainty.
Yōkai poses three key challenges to cooperative agents: First, achieving a high score requires agents to end a game as early as possible, resulting in a higher risk of failure. Second, ending the game early requires interpreting others’ actions and tracking their beliefs to form correct own beliefs over the colours of unobserved cards (see Figure 1). Third, agents must use hints efficiently as grounded communication, but a single misplaced hint can render the game unwinnable.
3 Preliminaries
We model Yōkai as a decentralised partially observable Markov decision process (Dec-POMDP) Oliehoek and Amato (2016), formally described as a tuple $\mathcal{M}=\langle\mathcal{N},\mathcal{S},\Omega,O,A,\mathcal{R},\mathcal{T},\gamma\rangle$. $\mathcal{N}$ is the set of agents where $i\in\mathcal{N}$, $1\leq i\leq|\mathcal{N}|$ denotes one specific agent and i-i denotes all other agents. Set $\mathcal{S}$ denotes the possible true states of the environment, with $s\in\mathcal{S}$ being a single state. At timestep $t$ an agent in state $s_{t}$ receives a partial observation $o^{i}_{t}\in\Omega$ through the observation function $O(s_{t},i)$. The action-observation history per agent is denoted by $\tau^{i}$. Following Sclar et al. Sclar et al. (2022), an agent with no ToM chooses an action $a^{i}\in A$ only based on their observation of the environment $\pi^{i}(a^{i}|o^{i}_{t})$. Agents with zeroth-order ToM incorporate memory, conditioning their policy on a hidden state $h^{i}_{t}$ in addition to their current observation $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t})$. Furthermore, agents that reason about the mental states of others estimate their hidden states $\hat{h}_{t}^{(j)}$ for $j\neq i$ and condition their action on all $\hat{h}_{t}^{(-i)}$: $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t},\hat{h}_{t}^{-i})$. Based on the joint action performed by all agents $\bm{a}=(a_{1},\dots,a_{n})$ the environment returns a shared reward $r_{t}=R(s_{t},\bm{a})$ and then produces a new state based on the state transition function $\mathcal{T}$. Finally, using the discount factor $\gamma^{t}$, the agents optimise the discounted return of a trajectory given by $R(\tau)=\sum_{t}\gamma^{t}r_{t}$.
4 The Yōkai Learning Environment
The Yōkai Learning Environment implements the game logic with two exceptions: First, we support configurable game sizes (e.g., a 3x3 version with nine cards and three colours). We refer to it as 3x3 while the original game is referred to as 4x4. The 3x3 two-player game has four hint cards: one one-colour hint and three two-colour hints. Second, our game is restricted to a $g\times g$ grid to efficiently compute legal moves. We pick $g$ large enough not to limit gameplay in practice and confirm this experimentally, see Appendix B. Our implementation is based on JaxMARL Rutherford et al. (2023) and is fully GPU-accelerated. It allows end-to-end learning on a single GPU at hundreds of thousands of steps-per-second, see Table 3.
Yōkai as a Graph
We opted to model Yōkai as a graph since it is a natural formulation of the relational nature of the environment and allows for an efficient implementation in JAX. At time $t$, the state is a graph $G_{t}(V,A)$ with cards and hints as vertices $V=Y\cup H$ and adjacency matrix $A$. Within $G_{t}$, cards are vertices $V=Y\cup H$ and connected to neighbours via edges represented by an adjacency matrix $A$. Cards and hints are identified by their index $0\leq i<|Y|$ or $0\leq j<|H|$. Hint cards are not connected to any card. For example, in the 3x3 version with nine cards and four hints, colours (e.g., blue, green, red) and hints are randomly assigned each episode, and the graph $A\in\{0,1\}^{13\times 13}$ is updated when cards move. With this setup, moving a card $i$ boils down to obtaining $A^{\prime}$ from $A$ via $A^{\prime}=\text{move}(A,i,n)$ where in move the row and column vector at index $i$ of $A$ are replaced with a binary encoding of the new neighbours $n$. Additionally, a configuration is legal if all cards are connected, computed via a reachability matrix over $A$. Determining whether the game has been won then boils down to computing
$\bigwedge\limits_{c\in C}\texttt{isConnected}(A_{c})$ (2)
where $A_{c}$ denotes the submatrix of $A$ obtained by deleting all column and row vectors not associated with colour $c$, i.e. by finding the set of indices $\{j|0\leq j<|Y|\land Y[j]\neq c\}$ where $Y[j]$ denotes accessing the tuple $Y$ at index $j$. To compute the state-dependent set of legal move actions $A(s)\subseteq A$, we run $\texttt{isConnected}(A^{\prime})$ on all $A^{\prime}$ obtainable by moving a card to a position in the $g\times g$ field.
Observation Space
Agents receive graph-based or image-like observations encoding card properties (e.g., colour, position, lock state) and hint features. Observations include both public (topology and hints) and private (card features) information, leading to asymmetric knowledge. Graph observations encode cards as an adjacency matrix $A_{t}\in\{0,1\}^{|Y|\times|Y|}$, and their properties as node features $F_{t}\in\mathbb{R}^{|Y|+|H|\times f}$, where $f$ is the feature dimension. In image-like observations, card features $F$ are embedded into a $g\times(g+1)\times f$ tensor based on their positions, while hint card features are embedded into column $g+1$. Agents observe which cards others inspect, but not their content. Partial observability forces agents to seek information and track shared knowledge.
Action Space
Actions in the YLE are represented by a categorical action space. Additional players increase the number of hints and, thereby, the number of possible actions. The action space includes observing pairs of cards, moving cards, placing or revealing hints, ending the game, and a no-op. The action space size is the sum of the number of cards to look at $|Y|$, the number of target positions any card can be moved to $N_{moves}=g^{2}|Y|$, and the number of hints $|H|$. The number of cards a hint can be placed onto $N_{hint\_moves}=|H|\cdot|Y|$, $1$ action for no operation and $1$ action to end the game. The exact size of the action space depends on the setup but ranges from $\sim 1,000$ to $\sim 3,000$ in practice, all details are in Appendix B.
Reward
The score in Equation 1 is only given when the game ends successfully, and sometimes can be negative due to hint misplacements. Agents, however, may learn that it is better to avoid grouping cards altogether to avoid penalties. To address this, we modify the reward function. If the game ends without a win, we give a small negative reward to discourage this strategy:
$R=\begin{cases}S&\text{if game won}\\ -d_{a}-(|C|-c_{a})-i_{h}&\text{otherwise.}\end{cases}$ (3)
where $c_{a}$ is the number of correctly formed clusters, $i_{h}$ is the number of incorrectly placed hints, and $d_{a}\in\{0,1\}$ indicates whether the game was ended early. Since rewards are sparse, we shaped only the training reward to accelerate learning. We additionally reward observing a previously unobserved card (+1+1), grouping a single colour, and placing a correct hint (both +1+1) or an incorrect hint (1-1).
5 Baselines, Experiments, and Results
To evaluate the spatio-temporal belief tracking capabilities of agents in the YLE, we explore five complementary research questions (RQs). We begin by comparing agents with different neural architectures to establish baselines, then progressively test their capacity for belief modelling, generalisation, and reasoning in increasingly complex scenarios:
RQ1
How well do different policies reason over their memory (zero-order ToM)?
RQ2
Does learning belief representations and explicit action encoding improve agents’ performance?
RQ3
Do agents learn belief‑tracking strategies that let them coordinate with new partners?
RQ4
Do agents represent beliefs in their internal activations?
RQ5
Can agents scale to higher-order ToM scenarios involving four players?
Setup
We use the two-player 3x3 YLE version, which already presents a significant challenge for RL agents. Agents are trained using MAPPO Yu et al. (2022) for $10^{9}$ steps and evaluated on 1,000 games, across three random seeds. The maximum episode length was $32$ for this setting. We report four metrics: return (R), success rate (SR), number of colours clustered (NC) and the rate of successfully ending early (SEE). SEE captures how often agents end the game early and still win – a decision that requires strong belief inference under uncertainty that comes with significant risk of failure. We tuned key hyperparameters for our models (e.g., encoder size, depth, etc.); Training details are in Appendix C.
5.1 RQ1a: Perfect memory is not sufficient to solve the YLE
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
Human - $5.7\pm 0.2$.5) $95.0\pm 2.5$.8) $3.0\pm 0.0$.0) $65.0\pm 4.8$.8)
GCN-M perfect $0.9\pm 0.0$.5) $26.8\pm 1.3$.4) $1.6\pm 0.0$.3) $0.0\pm 0.0$.0)
GATv2-M perfect $1.9\pm 0.0$.4) $58.2\pm 1.3$.5) $2.3\pm 0.0$.2) $0.0\pm 0.0$.0)
RM-M perfect $1.6\pm 0.1$.1) $50.7\pm 3.5$.0) $2.2\pm 0.0$.6) $0.0\pm 0.0$.0)
CNN-M perfect $2.3\pm 0.0$.3) $71.8\pm 0.8$.8) $2.6\pm 0.0$.1) $0.0\pm 0.0$.0)
CNN-LSTM implicit $0.2\pm 0.0$.1) $3.9\pm 0.4$.1) $0.8\pm 0.0$.3) $1.0\pm 0.1$.8)
CNN-GRU implicit $0.3\pm 0.0$.2) $2.3\pm 0.0$.3) $0.7\pm 0.0$.1) $2.1\pm 0.1$.8)
CNN-S5 implicit $0.1\pm 0.0$.1) $3.0\pm 0.3$.8) $0.7\pm 0.0$.2) $0.1\pm 0.0$.2)
CNN-TrXL explicit $1.5\pm 0.0$.3) $43.5\pm 0.9$.7) $2.1\pm 0.0$.2) $0.0\pm 0.0$.0)
ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
Table 1: Agent performance on the 3x3 YLE. The first block compares encoder architectures with access to perfect memory (-M). The second block evaluates different temporal architectures with the best-performing CNN encoder. There, we compare implicit and explicit memory modules. Best results are in bold. ToMTrack (final block) is the best method overall when not given perfect memory.
Figure 3: Training curves for the 3x3 YLE, with bands showing the standard error over three seeds. Left: comparison of encoders with perfect memory. Right: comparison of memory architectures.
To evaluate if agents can reason about their own memory (zero-order ToM), we first compared different encoder architectures in a perfect memory setting: In this setting, agents were given perfect memory by including all previously observed cards in their observations. This experiment also served to set a performance upper bound. Given the YLEs spatial reasoning task and graph-based nature, we compared CNN-, GCN- Kipf and Welling (2016), GATv2- Brody et al. (2021), and Relation Module (RM)-based encoders Zambaldi et al. (2018). Architectural details are provided in Appendix C.2. To anchor results, we additionally establish a human performance baseline based on results from 25 games played by 5 participants (ages 24 - 28; see Appendix F).
As shown in Table 1 (top), all agents severely underperform the human baseline. Among agents, despite YLEs graph-based nature, the CNN encoder outperformed both GNNs and the RM in all metrics except SEE (see Figure 3, left). We hypothesise this is due to YLE’s dynamic spatial layout: card positions and relationships change frequently, which may be better captured by CNNs than by message-passing. In contrast to humans (SEE 65%), agents failed to solve the task reliably or finish early. This suggests that the YLE requires more than memory: It demands the ability to reason about asymmetric knowledge to maintain common ground over time and to act under uncertainty.
5.2 RQ1b: Explicit memory modules show stronger zero-order ToM than implicit ones
While the perfect memory setting establishes an upper bound, it assumes agents retain all past observations without limitation. To better reflect realistic constraints and to asses whether agents can reason over their own internal knowledge representations, we evaluated models with different temporal architectures. This tests their capacity to track and use information over time – a key aspect of zero-order ToM. We tested LSTM- Hochreiter and Schmidhuber (1997), GRU Cho et al. (2014), S5- Smith et al. (2023), and TransformerXL-based (Dai et al., 2019, TrXL) architectures, paired with the best-performing CNN encoder. S5 has recently excelled on memory tasks in Samsami et al. (2024), while TrXL includes an explicit memory mechanism shown effective in memory-intensive RL Parisotto et al. (2020).
As shown in Table 1 (middle) and Figure 3 (right), TrXL outperformed all other temporal models ($p<0.05$, one-sided paired t-test). Nonetheless, there is still a performance gap between TrXL and the best perfect memory baseline CNN-M. Contrary to expectations from prior work Samsami et al. (2024), S5 offered no clear advantage over traditional RNNs. LSTM, GRU, and S5 agents occasionally ended games early. This behaviour appears to reflect a form of gambling, as agents terminate the game after very few actions and without sufficient information-gathering. This suggests a failure to track beliefs over time, as agents that end early do so without sufficient inference – contrasting with human players, whose early-ending behaviour reflects confident ToM reasoning. This suggests a failure to reliably recall past observations, leaving guessing as the primary strategy. This tendency is further reflected in the short average game length ($22.9\pm 15.5$, $14.0\pm 15.6$ and $30.4\pm 1.42$, respectively). TrXL agents never finish early.
Taken together, these results indicate that explicit memory architectures, such as TrXL, are better suited for zero-order ToM reasoning. Nonetheless, the gap in performance relative to the perfect memory baseline demonstrates the difficulty of learning robust belief-tracking strategies in YLE.
5.3 RQ2: Modelling beliefs over cards and distilling partner policies improves results
Figure 4: Our ToMTrack agent generates beliefs over card colours using an auxiliary prediction head.
Figure 5: Left: compares return. Right: shows success rate. ToMTrack trains at least $1.5\times$ faster.
Given the partial observability of the YLE, agents must maintain a belief $b$ over the hidden colours $c$ of all cards. To support this, we designed ToMTrack, a method for learning an implicit model over cards’ colour (see Figure 4). ToMTrack is an online representation learning RL approach inspired by previous work on offline belief representation learning Wang et al. (2023). At each timestep $s_{t}$, the policy $\pi$ should learn useful representations that contain information about the cards’ colours, no matter if they have been previously observed or inferred (first-order ToM, recall Figure 1). We thus want our policy to model $p(c_{i}|s_{t}),\,\forall\,0<i<|Y|$. To achieve this, we added an auxiliary belief classification head $f_{\text{belief}}$ to the policy network. This head maps the intermediate representation learned by the policy to a distribution over possible card colours. During training, we used the ground truth card colours as labels. We used an auxiliary belief prediction task Jaderberg et al. (2017) to train $f_{\text{belief}}$. Specifically, we added a cross-entropy loss $\mathcal{L}^{\text{CE}}$ to the actor, $\mathcal{L^{\prime}}_{\text{a}}=\mathcal{L}_{\text{a}}+\lambda\mathcal{L}_{a}^{\text{CE}}$, and critic loss, $\mathcal{L^{\prime}}_{\text{c}}=\mathcal{L}_{\text{c}}+\lambda\mathcal{L}^{\text{CE}}_{c}$; where $\lambda\geq 0$ is a hyperparameter.
In addition, since information is often implicitly communicated using actions, we add a second input encoder to explicitly process the last action taken by the active agent. This enables the policy to explicitly interpret observed behaviour and update its beliefs about card colours accordingly, see Figure 1.
We also evaluated other baselines based on the simplified action decoder (SAD) Hu and Foerster (2020) or intrinsic motivation, but found that they underperformed, see Appendix H. In contrast, ToMTrack outperforms CNN-TrXL ($48.3$ vs $43.5$ SR) and decreases the gap to CNN-M ($71.8$ SR) (Table 1). These results demonstrate that modelling latent beliefs and interpreting others’ actions improves agent coordination.
5.4 RQ3: Agents do not exhibit transferable models of partner belief inference
Figure 6: Zero-shot coordination performance measured by SR. We report additional metrics in Appendix E.2. Self-play agents do not learn to coordinate with others. Other-Play agents (OP) have some success at partner generalisation. However, the additional complexity of YLE requires more symmetries as input to OP (+C+R) compared to Hanabi, which only requires +C.
Figure 7: ToM reasoning influences Refer to caption ’s decision. Without ToM reasoning, no move is optimal. A zero-order agent (memory-based) may move either the red or blue card, as both would group two cards. A first-order agent that correctly interprets Refer to caption ’s action (see Figure 1) will recognise that moving the blue card completes a full cluster. Agents are rewarded during training for forming complete clusters. An example of second-order ToM reasoning can be found in Appendix B.3.
Effective strategies for tracking one’s own and others’ beliefs should generalise to novel partners. We test this via zero-shot coordination (Hu et al., 2020, ZSC), where agents must cooperate with independently trained partners at test time. Success in ZSC indicates that agents do not rely on partner-specific conventions but instead learn to track beliefs, interpret actions, and reason about others’ knowledge in a generalisable way.
We first evaluate ZSC performance of the previously trained CNN-M and ToMTrack agents using cross-play (XP), where agents are paired with separately trained partners. Figure 6 shows that XP SR drops by $88.5\%$ on average, indicating a strong reliance on learned conventions.
To address this, we retrain agents using other-play (Hu et al., 2020, OP), which prevents overfitting to arbitrary conventions by exploiting environment symmetries. In Hanabi Bard et al. (2020), colour symmetry alone (OP+C) suffices for strong ZSC Hu et al. (2020). We find that this is not the case in YLE. As shown in Figure 6, while both CNN-M and ToMTrack trained with OP+C improve XP performance, a significant SP-XP performance gap remains. Unlike Hanabi, YLE introduces spatial-temporal complexity and thus additional symmetry classes (see Appendix C.4), e.g. agents could form conventions around locations. We therefore also test OP+C+R, using rotational symmetries to break spatial conventions. This improves XP further for CNN-M and ToMTrack. For CNN-M it nearly closes the SP-XP gap ($33.2$ vs $27.4$).
However, while OP+C+R improves XP, this comes at a significant gap in overall SR (e.g. $86.5$ vs $41.5$ SP in CNN-M), suggesting that non-OP agents rely heavily on spatial and colour conventions. When these are unavailable, coordination suffers – implying that belief tracking and action interpretation remain difficult to learn directly. This suggests that CNN-M and ToMTrack do not acquire general belief strategies out-of-the-box, though OP-trained variants may begin to.
Agent Memory T0 $\downarrow$ T1 $\downarrow$ Wrong $\uparrow$
CNN-M-OP+C+R Perfect $13.8\pm 0.0$.5) $86.5\pm 1.4$.2) $30.3\pm 0.3$.1)
ToMTrack-OP+C+R Explicit $12.6\pm 0.0$.9) $98.8\pm 2.8$.3) $28.2\pm 0.2$.7)
Table 2: Results from the diagnostic ToM test in Figure 7. We report the average rank (based on action probabilities) of zero-order (T0), first-order (T1), and incorrect actions (e.g., mismatched card colours). We find that agents fail to make zero- or first-order inferences in this simple test.
To further test agents’ ability to perform first-order ToM reasoning, i. e, to infer a partner’s beliefs and intentions, we designed a diagnostic task (see Figure 7). In this test, only agents with first-order ToM will correctly identify that they can group all blue cards with one move, despite not being able to observe all blue cards first-hand. Grouping blue cards is optimal as it increases the number of card clusters by one, increasing the chance of finishing successfully or possibly early.
As shown in Table 2, we find that even the best-performing ZSC agents fail to select optimal actions and instead make moves consistent with either zero-order or no ToM reasoning. This shows that, despite OP improvements, agents lack transferable models of belief inference.
In summary, agents in YLE struggle to track beliefs of novel partners. Unlike Hanabi, success in YLE likely cannot be achieved by symmetry-breaking alone: as additional symmetries are removed, performance declines. This suggests the bottleneck is not symmetry identification but rather the absence of learning mechanisms that foster generalisable ToM reasoning.
Figure 8: Probing accuracies for each step. Memory Oracle observes new cards whenever possible and recalls them perfectly. Agents fail to track colours over time, especially after the mid-game.
Figure 9: Performance comparison of the two-player (ToMTrack) and four-player (ToMTrack-4P) 3x3 YLE. 4-player YLE requires tracking common ground between more players, and results in lower performance with higher variance.
5.5 RQ4: Agents struggle to represent knowledge of self and others over longer timespans
To succeed in the YLE, agents must maintain a belief over all cards’ colours. However, our results so far suggest that models are not using beliefs correctly. To identify whether agents fail to represent beliefs or instead only fail to use them, we apply linear probes Alain and Bengio (2016) to the agents’ hidden activations at each timestep of the game, to predict the colour of each card individually Matiisen et al. (2022); Bortoletto et al. (2024a, b); Zhu et al. (2024). This approach assesses whether belief-relevant information is linearly decodable from the learned embeddings – whether the card’s colour was directly observed, inferred, or remains uncertain (see Appendix D for details).
Figure 9 shows that agents start to form internal representations accurately in the early-to-mid-game. However, performance degrades in the later stages. This suggests that agents struggle to retain and update beliefs over longer timespans, especially as cards move and the spatial arrangement grows more complex. Together, this indicates a challenge in representing dynamic beliefs in YLE.
5.6 RQ5: Agents struggle to scale belief reasoning in four-player YLE
To test whether agents can scale to higher-order Theory of Mind (ToM) reasoning, we evaluate them in the four-player version of YLE. In this setting, each player observes fewer cards directly, and knowledge is distributed across more teammates. As a result, coordination increasingly depends on inferring what others know and what others believe about each other’s knowledge. For example, reasoning about whether agent 1 understands what agent 2 knows about agent 3’s interpretation of agent 4’s hint involves third-order ToM.
This demand is reflected in performance: as shown in Figure 9, ToMTrack exhibits increased variance in both SR and R, and mean SR drops by over $10$ percentage points compared to the two-player setting. Note that the four-player game allows higher absolute returns due to more hint cards (see Figure 9, left). However, as reflected in the success rate, this does not suggest better coordination. Our results suggest that scaling ToM reasoning in multi-agent RL remains an open challenge, especially in settings that require maintaining shared beliefs across multiple agents over time.
5.7 Key takeaways & Limitations
Across all RQs, agents, even with explicit memory or belief modelling, fail to maintain accurate beliefs or coordinate reliably over time. They overfit to partner-specific conventions and struggle with belief tracking over longer horizons. Notably, humans successfully end games early in 65% of cases, while agents very rarely do so – even with perfect memory or belief modelling (see Table 1). This finding positions SEE as a compelling metric for assessing belief reasoning, as it reflects agents’ ability to infer shared beliefs and identify coordination success. We present two main takeaways: First, ToM reasoning must be evaluated with novel partners to avoid mistaking brittle conventions for genuine ToM — ToM tests must acknowledge the generality constraint Rakoczy (2012). Second, YLE offers a challenging environment for assessing common ground reasoning under uncertainty, even in its simplest 2-player 3×3 setting. It supports future work to explore scaling to more agents and the original game size. Both enable an investigation of increasingly complex ToM settings.
While the YLE significantly expands our ability to evaluate key elements of ToM and advance cooperative agents, it also has limitations: It currently abstracts away modalities used in human ToM, such as language or gaze, and focuses on isolating belief-based reasoning. Additionally, it focuses solely on cooperation and does not address belief modelling in competitive or mixed-motive settings. These simplifications allow for focused evaluation of core ToM challenges before addressing the full complexity of human ToM in more real-world settings.
6 Related Work
Current benchmarks study ToM reasoning in AI mostly from a static observer’s perspective Baker et al. (2011, 2017); Rabinowitz et al. (2018); Bara et al. (2021, 2023); Fan et al. (2021); Duan et al. (2022); Kim et al. (2023); Bortoletto et al. (2024c); Gandhi et al. (2023), framed as a supervised learning task – an approach with notable limitations Aru et al. (2023); Bortoletto et al. (2024a). Other benchmarks evaluated agents in interactive environments, such as the Hanabi Learning Environment (HLE) Bard et al. (2020) and SymmToM Sclar et al. (2022). In contrast to Hanabi, Yōkai requires spatial-temporal belief reasoning, requiring agents to bind beliefs to moving cards Fernandez et al. (2023). Compared to SymmToM, YLE offers greater strategic depth and stronger pressure to develop ToM. In SymmToM, heuristics outperform all baselines and communication is unrestricted, while in YLE, misused hints can irreversibly lock cards in losing positions. Unlike Hanabi and SymmToM, YLE also introduces a high-stakes belief-based decision: agents may end the game early for a higher payoff, sacrificing future information and relying more on their beliefs.
A second line of work investigated whether neural networks represent mental states internally Bortoletto et al. (2024a); Zhu et al. (2024). For example, Matiisen et al. Matiisen et al. (2022) probed DRL agents for intention representations in cooperative navigation. We extend probing to YLE, where beliefs evolve dynamically and asymmetric knowledge emerges over time – showing the value of probing in a more complex environment.
Finally, YLE contributes to the broader family of cooperative multi-agent RL (MARL) benchmarks Samvelyan et al. (2019); Carroll et al. (2019); Bard et al. (2020); Ruhdorfer et al. (2024), but with a strong emphasis on ToM and common ground reasoning. While many techniques were developed for Hanabi, our results indicate they do not transfer directly to the YLE, thus highlighting the new challenges posed by our benchmark and requiring new modelling approaches.
7 Conclusion & Future Work
We introduced the Yōkai Learning Environment – a multi-agent RL benchmark where effective collaboration requires belief tracking over space and time. Implemented in JAX, YLE supports end-to-end GPU training at hundreds of thousands of steps per second. We show that RL agents on the YLE struggle to retain knowledge, to generalise across partners, or to identify when coordination has succeeded – a high-stakes decision reflected in their inability to end games early. YLE thus offers a scalable, diagnostic testbed for advancing common ground reasoning in collaborative AI.
While our work focused on evaluating RL agents, future work might use this benchmark to evaluate other systems, like language agents Xi et al. (2023); Wang et al. (2024). We envision that Yōkai can play a major role in isolating and assessing critical aspects of these systems that realte to memory, spatial reasoning and common ground tracking.
Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C. Ruhdorfer. We especially thank R. Yang for implementing a prototype of the environment as part of her master’s thesis, and A. Penzkofer and L. Shi for numerous insightful discussions.
Appendix A Appendix: Infrastructure & Tools
Servers and software
We ran our experiments on two different server systems. Both are running Ubuntu 24.04. One is equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs. The other is equipped with NVIDIA H100-NVL GPUs with 94GB of memory and AMD EPYC 9454 CPUs. All training runs are executed on a single GPU only. We trained our models using JAX Bradbury et al. [2018], Flax Heek et al. [2023] and Optax DeepMind et al. [2020]. In terms of tools, we did our analysis using NumPy Harris et al. [2020], Pandas Team [2020], SciPy Virtanen et al. [2020] and Matplotlib Hunter [2007]. Our TransformerXL implementation is based on the one provided by Hamon [2024]. Our single-file implementation of MAPPO is based on the one provided by JaxMARL Rutherford et al. [2023] with additional inspiration taken from Hamon [2024].
Accesability and reusability
The Yōkai Learning Environment (YLE) integrates directly into JaxMARL and implements its standard environment interface. As a result, any baseline method implementation of JaxMARL can be adapted to our environment quickly and with minimal effort. This makes our open-source environment easily accessible to future researchers who will benefit from our reference implementations and those provided by the wider community. It also enables inter-environment comparison, facilitating broader benchmarking across cooperative multi-agent tasks. The project is accessible under https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env.
Training times and compute requirements
Training times vary between server setup and agent. Our longest training agent is ToMTrack-OP+C+R since it needs to compute memory updates for the TransformerXL and the asymmetric domain randomisation used in Other-Play. Experiments with ToMTrack-OP+C+R take 8 hours to complete per seed on our H100 setup and around 24 hours on our V100 setup. The CNN-M baselines train quickest:  6 hours on the H100 server and  14 hours on a V100 server. All our experiments are runnable on a single V100 with 32GB VRAM.
Appendix B Fast, Scalable, and Parallel Yōkai in Graph-based JAX
We first describe in detail how our environment operates and then share additional details on reinforcement learning in the YLE.
B.1 JAX-based Implementation
We carefully engineer our environment to be end-to-end usable on a GPU directly and make it just-in-time compilable using JAX. This design enables both simulation and learning to occur entirely on the GPU, avoiding costly GPU–CPU communication and significantly improving efficiency. As a result, the environment supports training at hundreds of thousands of steps per second on modern hardware, enabling fast experimentation with relatively modest compute resources.
This efficiency and scalability stem primarily from how we represent and update the environment’s state using adjacency matrices $A_{t}$ at each timestep $s_{t}$. In the following, we describe two practical examples of how this is achieved.
Efficiently computing legal moves
We always mask illegal actions and thus need to compute the set of legal actions at every step. While identifying observable cards and placable hint cards is computationally inexpensive, computing legal next moves is by far the most costly computation our environment performs.
To determine which card moves are legal in a given state $s_{t}$, the environment must simulate each possible action $a\in A$ and assess whether the resulting state $s_{t+1}$ is valid. A state is considered legal if (1) all cards are connected into a single, fully-connected graph via their neighbours, and (2) no cards overlap spatially.
To make this efficient, we compute all potential next states in parallel:
$S_{t+1}=\{s_{t+1}|s_{t+1}=\mathcal{T}(s_{t},a)\forall a\in\mathcal{A}\}.$ (4)
Each state $s_{t+1}$ is associated with an adjacency matrix $A_{t}$ that reflects the card layout after a move. For each candidate action (i.e., moving a card $y\in Y$ to a new grid position $(x,y)$), we first verify that the position is within bounds and the card is movable.
We then check whether each resulting state is legal by evaluating its adjacency matrix. Specifically, we compute the reachability matrix:
$R=(I+A)^{|Y|-1}$ (5)
The state is legal if all entries in $R$ are positive, i.e., $R[i][j]>0$ for all $i,j$. This test, implemented as a parallel matrix power operation, allows us to efficiently filter legal moves at each step.
While designing environments for JAX requires careful consideration, these efforts potentially yield significant speedups that make RL research accessible to a broad range of researchers.
Efficiently executing moves
In our graph-based setup, executing environment transitions is highly efficient. Observing a card simply involves unmasking its colour in the node feature matrix $F_{t}$. To execute a move, we update the adjacency matrix $A_{t}$ to reflect the new spatial configuration. Specifically, when a card with index $i\in[0,|Y|)$ is moved, we first remove its current edges by setting:
$A_{t}[j,i]=0\text{ and }A_{t}[i,j]=0\;\;\forall j\in[0,|Y|)$ (6)
Next, we compute a binary vector $v$ such that $v_{k}=1$ if card $k$ is a spatial neighbour of the new position of card $i$, and 0 otherwise. We then update the adjacency matrix by setting:
[…]
Resulting Performance
Num Envs $n$ 512 1,024 2,048
2-player 3x3 139,107 274,669 559,251
2-player 4x4 135,662 144,636 148,269
Table 3: We evaluate steps-per-second (SPS) of our environment on a single Nvidia H100 by taking random actions in $n$ parallel environments. Our environment scales to hundreds of thousands of SPS.
We present how our environment performs in terms of steps-per-second across many parallel environments. All experiments were run by taking 1000 random consecutive actions in $n$ parallel environments. Table 3 shows results. Our environment scales similarly to other recently proposed JAX-based RL environments, e.g. Rutherford et al. [2023]; Nikulin et al. [2023]; Ruhdorfer et al. [2024]; Matthews et al. [2024].
B.2 Reinforcement Learning in the YLE
We want to continue discussing additional aspects of reinforcement learning in the YLE.
Playing Yōkai on a Grid
Figure 10: Examples of start, winning, and illegal card configurations for a Yōkai game with nine cards and three colours.
While in the original game, cards can be placed in any legal configuration, our environment limits the configuration to be in a predefined field of size $g\times g$ where $g=9$ in the 3x3 and $g=10$ in the 4x4 version. This enables the efficient computation of legal moves but prohibits some rare card configurations. While computing the exact number of disallowed configurations is infeasible, we find them to be extremely uncommon in practice.
For example, let’s assume a 2-player 3x3 game on a 9x9 grid. In such a setting, there a 4 hints and thus a total of 8 move actions for both players together. Cards are initially placed in a square in the centre of the board, leaving 3 empty spaces in the grid towards either side. In theory, thus, if agents agree to start placing cards in a row towards one of these sides, they would run out of space towards the border after 3 moves. In practice, this is seldom a suitable strategy and thus does not usually occur in games of Yōkai. This is because players want to group cards and are limited by the number of moves. They thus do not tend to start grouping cards in random directions, but rather group cards where they initially find a certain colour.
To verify that this is also true for our agents, we count the number of times when, in an evaluation game, an agent places a card on a border position. Note this is an overestimation as it does not account for the cases where agents place a card on the border but do not intend to place further cards in the same direction. We find that policies rarely to never use to border region of the board. Our best-performing policy ToMTrack uses the border $0.0\pm 0.0\%$ of the time when evaluated over all three seeds in $1000$ games each. We also discuss this in section 5.7.
The action space in detail
To observe two cards, agents first pick from $|Y|$ and then from $N_{cards}-1$ cards. Next, an agent moves a card by picking one of $N_{moves}$ actions, which can be understood to take card $0\leq i<|Y|$ and place it at $(x,y)\in g\times g$. Finally, the active agent reveals a hint from the hint pile or places a hint onto a card by picking from $|H|+|H|\cdot|Y|$ actions. Consequently, the maximum episode length is $8|H|$ while the minimum is $1$. The number of actions is therefore
$N_{a}=1+|Y|+N_{moves}+|H|+N_{hint\_moves}+1.$ (8)
For example, the action space is of size $1,068$ for 2-player 3x3 games and $2,322$ for the 4x4 version.
YLE is a symmetric ToM environment
Sclar et al. [2022] defines symmetric ToM environments as environments that (1) have a symmetric action space, (2) feature imperfect information, (3) allow the observation of others, and (4) require information-seeking behaviour. YLE features all four and thus qualifies. First, the action space is symmetric (see above). Second, information is imperfect since agents can never observe all cards first-hand and, for most of the game, only know a very limited subset of all cards. Third, it allows the observation of others both using explicitly encoded actions, as in ToMTrack or by observing their moves indirectly via their observations. Fourth, information needs to be acquired by observing cards first-hand or, second-hand, by interpreting moves and reasoning about knowledge. Agents that do not seek information and, for instance, repentantly observe the same card, will fail the game.
Number of Players, Hints and the Difficulty
4x4 3x3
Number of Players 2 3 4 2 3 4
Number of 1-colour hint cards 2 2 3 1 2 3
Number of 2-colours hint cards 3 4 4 3 3 3
Number of 3-colours hint cards 2 3 3 - - -
Sum 7 9 10 4 5 6
Table 4: Number of hint cards according to the number of players and the environment version. The 3x3 version features three colours and thus has no three-colour hints.
How many and which hint cards are available depend on the environment version and the number of players. For a 2-player game, there are 2 hints of one, 3 hints of two and 2 hints of three colours, i.e. 7 in total. Since there are 4 possible one-, 6 possible two- and 4 possible three-colour hints, players get a random subset of these hints at the start of the game per category. We give a full overview of all possible variations in Table 4. The number of hints directly influences the maximum number of turns in the game. This is because the game ends when all hints have been used. This means that games take longer if more players are added. Note, though, that this usually does not make the game easier. Instead, the number of cards one player can directly observe themself decreases with the number of players. For 4x4 YLE in the 2-player version, any agent can observe up to 14 of the 16 cars themselves (8 of 9 in 3x3 Yōkai). This shrinks to 10 of 16 for the 4x4 version and 6 of 9 for the 3x3 version when 4 players take part. Note how in the 3- and 4-player versions, agents are thus required to both work with less first-hand information and need to rely more on hints and interpreting the moves made by those around them. Not only this, but agents are tasked to do this tracking for more players as well. Thus, adding players to the YLE increases the difficulty substantially in two ways.
Additional Details on Legal Positions
Figure 11: Valid target positions for the card with the blue dot. Cards can only be moved so that all cards remain connected via their sides. Cards with an orange dot cannot be moved at all legally. Example 1 shows the initial position. All cards can be moved in the initial configuration. Example 2 stems from the early game. Examples 3 and 4 display configurations that might be encountered during the mid or late game.
We display several examples of legal target positions given a card to be moved in Figure 11. These only display a subset of all legal moves since the set of legal moves is different for each card on the board. For some cards, there are no legal moves whatsoever. This typically occurs when a card occupies a position in a narrow band that connects two clusters, such as in example 3. In example 3, there is no conceivable target position for both cards marked with an orange dot that would keep the cluster of cards connected, except their current position. Note that cards might be movable in the future again, i.e. if other cards are moved such that cards are freed again. In example 3, for instance, if the card with the blue dot is placed above any of the cards marked with an orange dot, both cards gain one additional legal target position, which makes them movable. Moving cards in a way that keeps option value (i.e. that keeps many cards moveable) is thus recommended. Observe example 4, for instance, here players placed cards in such a way that for most cards, very few target positions are still legal. This will usually highly restrict the agents’ ability to sort cards.
B.3 Second Order ToM Reasoning Example
Figure 12: A second-order ToM reasoning example over multiple timesteps in 2-player YLE. Refer to caption beliefs that Refer to caption beliefs that he (Refer to caption ) knew that cards 1 and 2 are of the same colour. Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground. In the future, they can now both observe and move different cards to finish early as they have one less card that needs to be observed.
Figure 12 displays an example for second-order ToM reasoning in 2-player Yokai. It shows an agent ([Uncaptioned image] ) making inferences about what they assume the other agents ([Uncaptioned image] ) knowledge is due to their own previous action. In the example [Uncaptioned image] infers the colour of the card due to the behaviour of the other agent. [Uncaptioned image] reasons that because of how [Uncaptioned image] played, [Uncaptioned image] must have interpreted his move correctly. [Uncaptioned image] thus reasons that [Uncaptioned image] thinks that [Uncaptioned image] knew that cards 1 and 2 have the same colour and must go together. This is the case, [Uncaptioned image] made this inference correctly, and card 3 indeed is also blue. This reasoning allows agents to require less first-hand observations to get an overview of the board, potentially enabling them to end early.
Note that an interesting strategic aspect of Yokai is the amount of shared information agents acquire together (both [Uncaptioned image] and [Uncaptioned image] choose to observe card 2) versus how much information they try to acquire asymmetrically to then use to hint knowledge to others via hints or actions (cards 1 and 3). Agents thus effectively control how big their shared knowledge is and when to strategically expand it. This has interesting strategic depth. Agents will differ in their strategic preference here, and developing strategies that are suitable for a wide range of partners is therefore a challenge.
This is just one illustrative example of ToM reasoning in YLE. Belief reasoning in the YLE can range from recalling previously observed cards (zero-order ToM), to tracking what another player knows (first-order), to inferring what one player believes about another’s beliefs, including their beliefs about oneself (second-order and beyond).
Note that such higher-order reasoning steps are even likelier to occur in settings involving more partner agents.
Appendix C Training Details
We share all the details of the training here, including all hyperparameters used in training and all details on the network architectures and how we tuned them. In general, we have tuned many aspects of our policy to arrive at the strongest agents we could for our experiments, see below. Finally, the number of parameters per model is shown in Table 6. We took extra care in making these equal to make our results comparable.
C.1 Reinforcement Learning Details
Description Value
Optimiser Adam Kingma and Ba [2015]
Adam $\beta_{1}$ $0.9$
Adam $\beta_{2}$ $0.999$
Adam $\epsilon$ $1\cdot 10^{-5}$
Learning Rate $\eta$ $3\cdot 10^{-4}$
Learning Rate Annealing Yes (linear)
Maximum Grad Norm $0.5$
# Simulators $1024$
Discount Rate $\gamma$ $0.99$
GAE $\lambda$ $0.95$
Entropy Coefficient $0.01$
Value Loss Coefficient $0.5$
# PPO Epochs 44
# PPO Minibatches 44
# PPO Steps $128$
PPO Value Loss Clipped
PPO Value Loss: Clip Value $0.2$
PPO Value Loss: Scale Clip Value No
Reward Shaping Yes (linearly annealed)
Table 5: Hyperparameters of the learning process.
Our work employs a standard MAPPO algorithm Yu et al. [2022] for all experiments. MAPPO is an extension of PPO Schulman et al. [2017] to the multi-agent setting. It separates the actor and the critic network. Since the actor and critic are separated and the critic is discarded after training, MAPPO allows giving the critic access to privileged information during training. This leads to the good performance of MAPPO in cooperative multi-agent tasks, compare Yu et al. [2022], which is why we choose it in our work. We present all parameters of the learning process in Table 5.
C.2 Neural Network Architectures
Model # Parameters
GCN-M 317,048
GATv2-M 317,560
RM-M 354,168
CNN-M 322,156
CNN-LSTM 470,252
CNN-GRU 437,484
CNN-S5 388,524
CNN-TrXL 400,940
ToMTrack 404423
ToMTrack-SAD 417863
Table 6: Number of parameters of all models used in this work.
Our work employs several encoder architectures. We present all the details on them below. In general, we aimed to make comparisons fair by giving all compared models a similar number of parameters, also see Table 6. The biggest advantage in terms of parameter size was RM-M for the spatial reasoning and CNN-LSTM for the temporal reasoning experiment, but neither performed the best.
C.2.1 General Policy Architecture
In general, in our work, no matter the encoder, the final policy network always has the same structure. That is given an embedding $e_{\text{encoder}}$ produced by an encoder network $f_{\text{encoder}}$ and possibly a memory architecture, we pass it to a stack of dense layers and either a value or action head. We always use $4$ dense layers before the heads with ReLU Fukushima [1975]; Nair and Hinton [2010] as the activation function between the layers and a hidden dimension of $128$. The layers are initialised using orthogonal initialisation with a scale of $\sqrt{2}$ and bias 0. We have determined the number of policy layers and their size using hyperparameter tuning. We compared $2$, 33 and $4$ layers with hidden dimension $32$, $64$ and $128$. We found that the largest configuration outperforms the others. Additionally, for all neural network architectures detailed below, we also varied and explored: The number of encoder blocks (2, 3 and 4), the hidden dimension of the encoder blocks, the aggregation function after the encoder (mean, sum, multiply and concatenate) and whether to anneal the dense rewards (False and True). Finally, the value head produces a singular scalar output and is initialised using orthogonal initialisation with a scale of $0.00$ instead of $\sqrt{2}$. We initialise the action head in the same way but it produces logits for all possible actions instead. The critic network is always parameterised the same as the actor network, but the input is different. Since we are using MAPPO our critic receives as input a so-called world state that combines the individual observations of all agents. This is possible in MAPPO due to centralised training and decentralised execution Foerster et al. [2016]. As the world state, we concatenate the individual observations along the feature dimension.
C.2.2 GCN, GATv2 and the pre-embedding of Nodes
Figure 13: A sketch of the GNN and pre-embedding architecture used in this work.
We use two graph neural networks in our work, one based on graph convolutions Kipf and Welling [2016] and one based on the attention mechanism Brody et al. [2021]. We pick these two as they stem from two different families of graph learning architectures and they thus are representative of a larger body of work. In our work, both methods observe the graph directly as described in the paper. There are two additional implementation details to note, though. First, both graph architectures observe an input graph that features a virtual node, similar to the hub-nodes of Abdessaied et al. [2024] or the [VNode] token of Ying et al. [2021]. A virtual node is an empty node that connects to all other nodes in the graph. It solely serves the purpose of reducing hop distances between nodes. The motivation for employing them in the YLE is that since cards can be distant from each other in terms of hop distance, information from one card might not propagate to the next during message passing. I.e. imagine the GNN observes two blue cards on opposite sides of the current card cluster. If the number of message-passing rounds is too low, this information might never be aggregated during message-passing. In RL, neural networks tend to be comparatively small, which is also true in our work. This also limits the number of message-passing steps to be small (2 in our case). With a virtual node, however, this is not an issue since the virtual node aggregates information from all other nodes in one single message-passing step. Second, our work uses pre-embeddings of node-feature modalities as described in Bortoletto et al. [2024d]. We take the node features of size $n\times f$ and embed them per modality via a separate dense encoder before then concatenating all of the resulting vectors into new node features shaped $n\times f^{\prime}$. $f^{\prime}$ depends on the size of these dense layers. We sketch the overall approach in Figure 13. In general, we found that embedding the node features per modality before using them in the GNN improved performance following earlier work, i.e. see Bortoletto et al. [2024d]. Specifically, for each modality $m\in M$ we train a dense layer $f_{m}$ that embeds the modality per node into an embedding $e_{m}\in\mathbb{R}^{b\times t\times n\times 2}$ for the current player embedding or $e_{m}\in\mathbb{R}^{b\times t\times n\times 4}$ for all other modalities where $b$ is the batch size, $t$ the time and $n$ the number of nodes. All embeddings are then concatenated into a node embedding $e_{\text{node}}=e_{\text{colour}}\oplus e_{\text{position}}\oplus e_{\text{locked}}\oplus e_{\text{seen}}\oplus e_{\text{id}}\oplus e_{\text{is\_current\_player}}\oplus e_{\text{substep}}$ with $e_{\text{node}}\in\mathbb{R}^{b\times t\times n\times f^{\prime}}$ with $f^{\prime}=26$ using the configuration described. We then add self-connections, the virtual node using a zero-vector $e_{\text{virt}}\in\{0\}^{f^{\prime}}$ and connect the virtual node to all real nodes before passing them to our graph encoder $f_{\text{GNN}}$. The output dimension of the graph encoder is $64$. For GATv2, we use $4$ heads. We apply $4$ graph layers in the relational encoder comparison. The resulting embedding is then concatenated along the node dimension and fed to the rest of the policy.
C.2.3 Relation Module
Our work adapts the relation module of Zambaldi et al. [2018]. We only use the self-attention mechanism of their proposed relation module and skip the CNN encoder they propose alongside the self-attention layers. This is because our setting deals with an environment where agents observe the position of cards directly as opposed to the setting of Zambaldi et al. [2018] that deals with image inputs directly. Thus, in our relation module, non-local interactions between all cards are simply computed using self-attention directly. Similarly to the GNN encoders, we also pre-embed the modalities of all nodes using the same mechanism as discussed above. The resulting embedding is then also passed to the remaining policy. Our relational module is also applied $4$ times, features $4$ heads, and its output dimension per layer is $64$.
C.2.4 CNN
Our work compares a CNN encoder and determines it to be the best-performing model. As described earlier, the YLE returns an image-like observation to be processed by our CNN. The exact size of the input tensor to the CNN depends on the YLE settings, but in 2-player 3x3 YLE, the input dimension is $9\times 10\times 13$, which we will use as a running example. It includes the same features as discussed above, with some minor changes. The first three channels include a one-hot encoding of cards’ colours, the next three an encoding of hint colours, followed by two channels for the position, one for whether the card is locked, one for whether the observing agent is the current player, one for the number of the current substep (1-4) and one that identifies the card via an identifier number. Our convolutional stack uses $4$ 2d convolutions with $64$ filters each and ReLU applied between them. The kernel size is $(3,3)$, we apply a stride $1$ and valid padding. The resulting embedding is flattened to produce the final embedding to be passed on to later layers.
C.2.5 LSTM and GRU
Next to the encoder architectures, our work compares several memory architectures. Here, memory architecture is used as a term to denote classical RNNs as well as structured state space or transformer Vaswani et al. [2017] architectures. LSTMs Hochreiter and Schmidhuber [1997] and GRUs Cho et al. [2014] are commonly used in RL and such are natural first approaches for our environment. They provide an implicit memory through their hidden state that is computed at every step from the current observation encoding and the previous hidden state. In both cases, we take the embedding generated by the encoder, $e_{\text{encoder}}$, pass it through a dense layer $f_{\text{pre\_rnn}}$ before finally applying either the LSTM or GRU on the resulting embedding. $f_{\text{pre\_rnn}}$ embeds $e_{\text{encoder}}$ to the hidden dimension of the RNN. As the hidden dimension of the RNN we pick $128$ in our work. We also experimented with using $256$ but found no difference in performance. We use orthogonal initialisation for both the GRU and the LSTM.
C.2.6 S5
As already described, we use S5 Smith et al. [2023] layers because of the great performance of structured state-space models on memory tasks in deep model-based RL Samsami et al. [2024]. Similar to LSTM and GRU, S5 only provides an implicit memory mechanism. We use a similar setup as with both the LSTM and GRU layers, i.e. we also use a $f_{\text{pre\_s5}}$ embedding layer that embeds $e_{\text{encoder}}$ to the hidden dimension of our S5 layer. We use $2$ S5 blocks, $1$ layer, place the LayerNorm Ba et al. [2016] before the S5 layer and use the activation function proposed in the original work Smith et al. [2023]. As before we use $128$ as the hidden dimension of S5.
C.2.7 TransformerXL
Finally, as already introduced TransformerXL Dai et al. [2019] is an explicit memory mechanism that shows great performance in memory-intensive RL tasks when used correctly Parisotto et al. [2020]. Our usage strictly follows the one described by Parisotto et al. [2020] as implemented by Hamon [2024]. Similar to the other memory architectures, we also first pass the embedding received by the encoder through a dense layer $f_{\text{pre\_trxl}}$ before handing the resulting embedding to the transformer. While LSTM, GRU and S5 use $128$ as their hidden dimension, we only use a hidden dimension of $64$ for both the dense layers within the transformer as well as the projection layers in multi-head attention. This is to keep the number of parameters consistent between the architectures. We furthermore configure the transformer to have $8$ heads and to use gating with a gating bias of $2$. We use a single transformer layer. Finally, we set the length memory window to $32$. $32$ is long enough to recall all of the previous moves in the 3x3 YLE. Memories from a past game – as marked by the done signal – are masked. We apply standard positional encoding based on sinus waves.
C.3 ToMTrack
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
ToMTrack w/o $f_{\text{belief}}$ explicit $0.9\pm 0.0$.5) $31.2\pm 1.4$.8) $1.8\pm 0.0$.4) $0.0\pm 0.0$.0)
ToMTrack w/o action explicit $1.4\pm 0.0$.4) $41.2\pm 1.1$.7) $2.0\pm 0.0$.3) $0.0\pm 0.0$.0)
ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
Table 7: ToMTrack ablations. We evaluate ToMTrack without the additional belief representation learning and the additional action encoding and find that only both together improve performance.
ToMTrack is a direct extension of the CNN-TrXL baselines and shares all the parameters mentioned earlier with both the CNN as well as the TransformerXL work. There are two extensions we add to arrive at ToMTrack. The first one – the auxiliary belief prediction task – is already mostly described in the main text. To generate the distribution over possible card colours, we add a colour prediction head $f_{\text{belief}}$ to our policy. $f_{\text{belief}}$ takes as input the embedding generated by the stack of dense layers and produces an output $e_{\text{belief}}\in\mathbb{R}^{b\times t\times|Y|c}$ that resembles the prediction over all cards and colours. For the 3x3 version $|Y|c=9\cdot 3=27$. To obtain the final predictions, we reshape $e_{\text{belief}}$ to be of shape $b\times t\times|Y|\times c$ and apply the Softmax function of the last dimension to get a prediction per card. The method only has one hyperparameter $\lambda$ – which balances the losses - and we pick $\lambda=0.1$ in this work. The second extension – the active action decoding – is realised via an additional input stream. Before passing the previous action to this stream, we first decode it into a reduced dimension that serves as a dense representation of that action of shape $4+|Y|+2+|H|+1$. The first $4$ bits encode the move type alias the substep. The next $|Y|$ bits encode the Yōkai card affected by the current action, if no card is affected, all are set to 0. The next $2$ bits encode the position affected by the action. For instance, if the previous active action was a move action, the resulting coordinates are encoded there. The next $|H|$ bits encode whether and if so which hint card is used in an action, and finally, the last bit encodes whether the previous action was to terminate the game. For a 2-player 3x3 YLE game $4+|Y|+2+|H|+1=20$. Note that this is a much denser representation compared to just encoding the raw number. The resulting encoding $e_{\text{action}}$ is processed by its own dense layer and then concatenated with the output of the encoder before being passed to the rest of the policy network. Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
C.4 Other-Play in Yōkai
Figure 14: An example of how both agents act in rotated and recoloured environments to break symmetries for zero-shot coordination.
The core idea of Other-Play (OP) is that during self-play training, agents learn to break symmetries of the environment in arbitrary ways that do not generalise to novel partners. In Hanabi, for instance, agents might form conventions around certain colours. OP presents agents with a special asymmetric domain randomisation that makes it impossible for agents to coordinate on breaking symmetries in certain ways. In Hanabi, this is achieved by a recolouring operation that shows each agent a different recolouring of the game state Hu et al. [2020].
Therefore, to train agents via OP for the YLE, one first needs to find all classes of symmetries that exist in an environment, and then, second, find an appropriate domain randomisation operation. Generally, since agents also observe coloured cards in the YLE, we can apply the same recolouring operation as Hu et al.. However, since the YLE features a spatial grid, additional symmetries exist around the location of cards. Even if agents can not coordinate on specific colours when we apply recolouring, they are, for example, able to coordinate on sorting the first colour they find in a specific place. Agents, for instance, might choose to observe the same two cards at the start and place one of them in the top left corner. This is an arbitrary convention (top left vs bottom right etc.). To prevent this, we apply asymmetric rotation to the agents’ observations. This way, when one agent places a card next to the top-left corner in their environment, their partner agents observe this as placing the card next to potentially any corner in different episodes. Training with rotated observations requires to also requires rotating the actions of the agent appropriately. We show an example of this process in Figure 14.
Note that these are not the only ways for agents to learn to break symmetries. Recent work Treutlein et al. [2021] showed that agents also might coordinate to break symmetries in arbitrary ways over multiple timesteps. It is impossible to list all possible classes of symmetries for an environment with enough complexity. This is a drawback of the OP algorithm as it requires them as input. In our work, we thus focus on showing that while Hanabi only requires breaking one symmetry for efficient ZSC while the YLE – due to its additional complexities – requires additional effort to improve ZSC performance, which is still far from optimal.
Appendix D Additional Probing Details
Over the duration of $1000$ games, we collect a probing dataset $\mathcal{D}=\{x^{(i)},y^{(i)}\}$. The probing datasets map the activations of an agent to a label $z$. A probe then is a function $g:f_{l}(x)\to\hat{z}$ and trained on $\mathcal{D}$ where $f$ is a neural network and $f_{l}(x)$ is the intermediate representation of $x$ at layer $l$. In our case, we train simple logistic regressions. We always probe before the action head of our agent. As highlighted in the main text, during all experiments, we train one and evaluate it on a hold-out test set $\mathcal{D}_{\text{test}}\subset\mathcal{D}$. Additionally, note that many of these tasks feature imbalanced datasets since the does-not-know label is far more likely to occur than any other label. We correct for that by weighing the loss per class and subsampling the does-not-know label to the largest other class. During all experiments, we train one probe per card – 9 in total – and average the scores computed on a hold-out test over all 9.
Appendix E Additional Experimental Results
E.1 Additional Training Curves
Figure 15: Return, success rate and numbers of clusters over time per encoder while given access to perfect memory.
Figure 16: Return, success rate and numbers of clusters over time per memory architecture.
Figure 17: Return, success rate and numbers of clusters over time for ToMTrack, ToMTrack-SAD, CNN-TrXL and the CNN-M architecture.
For completeness, we show how the most important metrics of our agents developed over time. Figure 15 shows the curves of the spatial-reasoning experiment, Figure 16 of the memory-reasoning experiment and lastly Figure 17 reports the same for ToMTrack.
E.2 Additional Zero-Shot Coordination Metrics
Figure 18: Zero-shot coordination performance measured by success-rate, average return and number of cards clustered. SP and OP+C are consistently outperformed by OP+C+R, showing that beyond cards’ colour, agents need to focus on other reasoning types.
Figure 19: We show training curve metrics for our agents trained with other-play to demonstrate convergence. Recall that these metrics measure self-play performance rather than ZSC. ZSC metrics are shown in Figure 6 and Figure 18.
In Figure 18 we show additional metrics for the ZSC evaluation in our main body of work. Training curves for our other-play agents are shown in Figure 19 to show convergence behaviour.
E.3 2- versus 4-player Games
Figure 20: ToMTrack in the two- and four-player 3x3 YLE. In four-player YLE more hints are available to be placed, and as such, we observe higher total rewards collected by the ToMTrack-4P agent, even though it wins fewer games and clusters fewer colours.
In Figure 20 we show how two- and four-player settings compare for our best model, ToMTrack.
Appendix F Human Performance in Yōkai
We collected human game performance in a setting that adheres to the same setting that artificial agents face, i.e. participants played with 9 cards on a 9x9 grid in a two-player setting. Humans played in person. A pair of subjects was instructed on the rules verbally. Each pair of subjects was given a single training round to make sure that all participants were familiar with the rules. To make results as comparable as possible, humans were instructed to avoid any form of non-verbal communication. The study was conducted under an umbrella agreement of the relevant institutional review board. Compensation was also handled under the rules of the relevant institution.
Appendix G Example Games
Figure 21: We show an example game in which ToMTrack wins in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. In turn 7, the agent finishes the final cluster of cards by correctly inferring that the card under the rbrb-hint must be red and not blue. In turn 8, agent 1 should have finished early for more reward, but did not recognise that fact and instead used the last hint card, thus finishing the game. Note that the agents tend to play quite conservatively. In turns 1 and 3, for instance, it gains knowledge of two separate green cards but only groups them in turn 5 after agent 1 places a corresponding gbgb-hint in turn 4.
Figure 22: We show an example game in which ToMTrack loses in self-play. For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom. We also show how many card clusters exist at this point per turn. Hint cards are highlighted by a light-blue border. Note that agent 0 in turn 1 finds a blue group early and manages to form a cluster quickly. Agent 0 confirms this by turn 3. In turn 6, the game is still winnable. However, because both agents over-focused on establishing a common ground around the blue cluster, they fail to realise this and in turn 7, agent 0 then blunders and misplaces the green card. From there, the game is lost under all circumstances. Additionally, note that in turns 4 and 6, agent 1 establishes the identity of all red cards but fails to recall that in turn 6, therefore also misplaying.
We show an example game in which ToMTrack wins in self-play in Figure 21 and one where it loses in Figure 22.
Appendix H Details on Other Attempts
While ToMTrack provides a mechanism for learning representations that represent card colours, it does not necessarily encourage the agents to maximise their shared or individual knowledge. Consequently, we were inspired to also experiment with a variant of ToMTrack that is not trained via an auxiliary task but rather uses $\mathcal{L_{\text{CE}}}$ as intrinsic motivation for both agents; ToMTrack-I. ToMTrack-I agents are encouraged to form an implicit belief model that aligns well with the true state of the environment by adding an intrinsic reward $r_{i}$ to the reward based on the negative cross-entropy loss of the actor classification head:
$r=r+\lambda_{\text{s}}*r_{\text{s}}+r_{\text{i}}\text{, where }r_{\text{i}}=-\mathcal{L}^{\text{CE}}_{a}.$ (9)
In this formulation, the agents thereby receive a higher reward when the belief model’s prediction loss $\mathcal{L}^{\text{CE}}_{a}$ is low. This approach is loosely inspired by Oguntola et al. [2023]. We originally assumed that this would achieve superior results but found that the intrinsic reward confused the agents to a point where they barely learned a useful policy altogether, falling below baseline performance (i.e. CNN-TrXL). Future work should study more sophisticated ideas based on this concept.
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
ToMTrack (SAD) explicit $0.7\pm 0.0$.4) $21.5\pm 1.3$.4) $1.5\pm 0.0$.4) $0.0\pm 0.0$.0)
Table 8: Additional ToMTrack-SAD results. This approach underperforms our baselines from Table 1 considerably and was thus excluded from the main paper.
Previous work on Hanabi has proposed the simplified action decoder (SAD) Hu and Foerster [2020]. SAD agents can take two actions during training that other agents can observe: a normal environment action and a greedy action. Other agents are allowed to observe both. The intuition is that during exploration, the true intentions of agents are blurred with explorative actions and thus less informative for others. Hu and Foerster [2020] implements their idea on top of epsilon-greedy Q-learning. Since during evaluation $\epsilon=0$, all actions are greedy and are given as input to both action inputs heads of other agents. During evaluation, $\epsilon=0$ and both actions are greedy. SAD showed improved performance in Hanabi, including the Bayesian action decoder Foerster et al. [2019]. We have adapted this idea as ToMTrack-SAD to our policy learning approach. Since we do not have an $\epsilon$ to set, we adopt their idea by also allowing agents to take two actions during training: One sampled from the agents’ produced action distribution is used as the environment action that gets executed, and one greedy action is determined as the legal action with the highest associated logit. The true environment action is duplicated during evaluation with similar reasoning.
ToMTrack-SAD is our extension of ToMTrack. In ToMTrack-SAD both actions are encoded as described earlier and concatenated before being fed to the same dense layer. Since we use MAPPO and no Q-learning approach as in the original work Hu and Foerster [2020], our agents pick their normal environment action given their current set of parameters $\theta_{a}$ in state $s_{t}$ via
$a_{\text{env}}\sim\pi(\cdot|s_{t},\theta_{a})$ (10)
and their second greedy action as
$a_{\text{greedy}}=\arg\max_{i}\pi(a_{i}|s_{t},\theta_{a}).$ (11)
Both are encoded and processed as described above when discussing ToMTrack. Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
Results in Table 8 show that ToMTrack-SAD performs worse than CNN-TrXL or other baselines. We speculate that including greedy actions as input is uninformative in an environment with many actions like the YLE.

### v2

@@
 Abstract
-Developing collaborative AI hinges on Theory of Mind (ToM) – the ability to reason about the beliefs of others to build and maintain common ground.
-Existing ToM benchmarks, however, are restricted to passive observer settings or lack an assessment of how agents establish and maintain common ground over time.
-To address these gaps, we introduce the Yōkai Learning Environment (YLE) – a multi-agent reinforcement learning (RL) environment based on the cooperative card game Yōkai.
-In the YLE, agents take turns peeking at hidden cards and moving them to form clusters based on colour.
-Success requires tracking evolving beliefs, remembering past observations, using hints as grounded communication, and maintaining common ground with teammates.
-Our evaluation yields two key findings: First, current RL agents struggle to solve the YLE, even when given access to perfect memory.
-Second, while belief modelling improves performance, agents are still unable to effectively generalise to unseen partners or form accurate beliefs over longer games, exposing a reliance on brittle conventions rather than robust belief tracking.
-We use the YLE to investigate research questions in belief modelling, memory, partner generalisation, and scaling to higher-order ToM.
+The ability to cooperate with unknown partners is a central challenge in cooperative AI and widely studied in the form of zero-shot coordination (ZSC), which evaluates an algorithm by measuring the performance of independently trained agents when paired.
+The Hanabi Learning Environment (HLE) has become the dominant benchmark for ZSC, but recent work has achieved near-perfect inter-seed cross-play performance, limiting its ability to track algorithmic progress.
+We introduce the Yōkai Learning Environment (YLE) – an open-source multi-agent RL benchmark in which effective collaboration requires building common ground by tracking and updating beliefs over moving cards, reasoning under ambiguous hints, and deciding when to terminate the game based on inferred shared knowledge – features absent in the HLE, where beliefs are tied to hand slots and hints are truthful by rule.
+We evaluate the leading ZSC methods, including High-Entropy IPPO, Other-Play, and Off-Belief Learning, which achieve near-perfect inter-seed cross-play in the HLE, and show that in the YLE they exhibit persistent SP–XP gaps, degraded early-ending calibration, and weaker belief representations in cross-play, indicating failure to maintain consistent internal models with unseen partners.
+Methods that perform best in the HLE do not perform best in the YLE, indicating that progress measured on a single benchmark may not generalise.
+Together, these results establish YLE as a challenging new ZSC benchmark.
 1 Introduction
-Effective collaboration among partners requires common ground – the shared knowledge, beliefs, and assumptions that enable them to interpret each other’s actions and respond appropriately Clark (1996).
-When common ground cannot be reliably maintained, collaboration often breaks down Klein et al.
-(2005).
-To establish common ground, humans use Theory of Mind (ToM) Premack and Woodruff (1978) to infer others’ mental states.
-Consequently, enabling computational agents to perform ToM reasoning has been identified as a critical step toward effective human-AI collaboration Klien et al.
-(2004); Dafoe et al.
-(2020, 2021).
-Research in cognitive science suggests that human ToM integrates multiple reasoning processes – including memory, spatial navigation, and mentalising Suddendorf and Corballis (2007); Mitchell (2009).
-Prior work in computational ToM has focused on the latter, for example, by predicting others’ behaviour from an omniscient observer’s perspective Baker et al.
-(2011); Rabinowitz et al.
-(2018); Gandhi et al.
-(2021).
-However, humans are rarely passive observers of a scene where they solely predict behaviour Sclar et al.
-(2022).
-Instead, they are active participants and must adjust their ToM reasoning to the interactive setting in which their own actions change the environment, influence the behaviour of others, and vice versa.
-This interactive structure necessitates tracking how common ground – the mutually recognised shared experience built through interaction Tomasello (2010) – evolves during collaboration.
-The goal of this paper is thus to
-propose an environment in which effective collaboration between computational agents requires building common ground by tracking beliefs over space and time.
-Figure 1: The YLE poses a challenging ToM reasoning task for cooperative agents in a dynamic environment.
+Effective collaboration among partners requires common ground – the shared knowledge, beliefs, and assumptions that enable them to interpret each other’s actions and respond appropriately (Clark, 1996).
+Maintaining common ground requires reasoning about what a partner knows, believes, and intends – commonly referred to as Theory of Mind (ToM) (Premack & Woodruff, 1978).
+Improving ToM reasoning in AI agents has emerged as an important requirement toward effective human-AI collaboration (Klien et al., 2004; Dafoe et al., 2020; 2021).
+A necessary precondition for human-AI collaboration is zero-shot coordination (ZSC) (Hu et al., 2020) which focuses on designing algorithms that, when run independently multiple times, produce compatible policies.
+The Hanabi Learning Environment (HLE) (Bard et al., 2020) is arguably the most widely used benchmark for ZSC and ToM (Hu et al., 2020; 2021; Fuchs et al., 2021; Cui et al., 2021; 2022; Muglich et al., 2022; 2025; Lauffer et al., 2025).
+However, close to perfect inter-seed cross-play performance has recently been reached (Sudhakar et al., 2025; Forkel et al., 2026), limiting its utility as a stress test for further ZSC progress.
+Consequently, the field would benefit from new benchmarks that can be used to track algorithmic advances without risking overfitting to the HLE.
+Figure 1: The YLE poses a challenging ToM reasoning task for ZSC.
 Agents cannot observe all cards in a single game.
-Successful play requires agents to remember and reason about others and their own mental states.
+Successful play requires agents to reason about other agents’ knowledge and beliefs.
 In the example, it’s Refer to caption ’s turn.
@@
 Refer to caption recalls from earlier that card 1 is blue.
-When Refer to caption moves card 1 next to card 2, Refer to caption can infer that the second card is likely blue as well.
-Refer to caption and Refer to caption can now assume this is common ground going forward.
-This is the Yōkai Learning Environment (YLE) [^1] Code available at https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env.
-– a novel multi-agent reinforcement learning (RL) environment based on the collaborative card game Yōkai.
+When Refer to caption moves card 1 next to card 2, Refer to caption can infer that the second card is blue as well.
+Refer to caption and Refer to caption can use this common ground going forward.
+Our solution is the Yōkai Learning Environment (YLE)– a novel multi-agent reinforcement learning (RL) benchmark for ZSC based on the collaborative card game Yōkai.
 In the YLE, players must collaborate to group face-down cards by colour over multiple rounds without direct communication (see Figure 1, left).
-Each round, one player privately observes the colour of two cards, moves one of them to a new position on the table, and may place a hint card to communicate information about unobserved card colours to others (see Figure 2).
-Importantly, players may end the game early for a higher reward, but at the risk of losing the game.
-This makes it a high-stakes, all-or-nothing decision that requires strong ToM reasoning to infer card colours from partners’ actions rather than first-hand observation.
-The design of YLE creates a uniquely challenging coordination task.
-Agents must: (1) remember the colours of moving cards to avoid redundant checks, (2) infer the beliefs and knowledge of others to establish and maintain common ground, (3) interpret teammates’ actions and hints as implicit belief updates, (4) use hint cards as grounded communication, and (5) assess when to end early based on inferred shared knowledge.
-These challenges require belief tracking over both space (shifting card locations) and time (multi-round gameplay with limited observability), and together reflect key components of ToM Premack and Woodruff (1978); Clark (1996); Pillow (1989); Southgate et al.
-(2007); Tomasello (2010) in a dynamic, interactive setting.
-Effectively establishing common ground requires aligning on ad hoc conventions for grouping cards with partners.
-Prior environments, such as Hanabi Bard et al.
-(2020), were used to study zero-shot coordination and ToM, but lack the spatial-temporal structure, evolving belief dynamics, and the challenge of ending early based on inferred knowledge.
-We evaluate a range of agents on YLE and highlight key limitations in their ability to track common ground in this novel spatial-temporal belief reasoning task.
-Our experiments show that even with access to perfect memory, agents struggle to collaborate effectively.
-Among agents without perfect memory, those with explicit memory modules perform best, and performance drops further in four-player settings.
-While belief representation learning improves performance, agents fail to generalise to new partners or maintain accurate beliefs over longer time horizons.
-Instead, they rely on brittle conventions rather than robust and generalisable belief tracking.
-Our contributions are:
+Each round, one player privately observes the colour of two cards, moves one of them to a new position, and may place a hint card to communicate information about card colours to others (see Figure 2).
+Hints in YLE can be multi-coloured and thus can be ambiguous.
+Furthermore hints can be placed on any of the yet unhinted board cards, including ones whose colour is not on the hint card.
+Hints are thus not required to be truthful.
+Crucially, players may choose to end the game early for a substantially higher return than completing the game conservatively, but at the risk of receiving a return of zero.
+Because most of the achievable return lies in successfully ending as early as possible, this creates a high-stakes trade-off: players must commit early to maximise reward, yet doing so requires sufficiently accurate and well-calibrated beliefs about the evolving board state.
+Premature ending leads to catastrophic failure, while excessive caution yields suboptimal return.
+To succeed in Yōkai, players must maintain accurate beliefs about dynamically moving cards, align interpretations of ambiguous hints and actions, and determine when evidence justifies early ending.
+The design of YLE therefore creates a coordination problem under sparse, ambiguously grounded information and high-stakes commitment.
+Unlike Hanabi, where agents observe their partners’ hands through a small number of fixed card slots and hints are truthful by rule, YLE reveals information locally in space: no player ever observes the full board, hints may be multi-coloured and need not be truthful, and cards move across the grid.
+As a result, beliefs must be repeatedly rebound to dynamically moving cards while trying to end early under uncertainty about the true board configuration.
+We show that ZSC algorithms, which achieve near-perfect inter-seed cross-play in HLE, exhibit persistent gaps between self-play (SP) and cross-play (XP) when evaluated in YLE, indicating coordination failures.
+We thus posit that (1) YLE is a timely new benchmark, as approaches that saturate HLE do not saturate YLE, and (2) YLE highlights limitations of existing ZSC methods that have so far been validated primarily in Hanabi.
+Our specific contributions are:
   1. 1.
-    A fast JAX-based YLE – a new cooperative multi-agent RL benchmark for belief tracking, common ground, and ToM, training at hundreds of thousands of steps per second.
+    We introduce the Yōkai Learning Environment (YLE) as an open-source JAX-based benchmark for zero-shot coordination in which agents must build and maintain common ground through spatio-temporal belief tracking, ambiguous communication, and high-stakes early termination.
   2. 2.
-    An empirical analysis showing that agents fail to coordinate effectively in YLE, and in particular fail to end games early – a behaviour humans achieve reliably.
+    We provide comprehensive empirical evaluations of leading ZSC algorithms that achieve near-perfect inter-seed cross-play in Hanabi – including Other-Play (Hu et al., 2020), Off-Belief Learning (OBL) (Hu et al., 2021), and High-Entropy IPPO (Forkel et al., 2026).
+In the YLE, these methods exhibit persistent SP–XP gaps, degraded early-ending calibration, and reduced coordination.
+To our knowledge, this is the first implementation of OBL outside Hanabi.
   3. 3.
-    A demonstration that belief representation learning improves performance, but agents still fail to generalise to novel partners or maintain accurate beliefs over time.
-  4. 4.
-    An evaluation of YLE in a four-player setting, revealing the increased difficulty of maintaining common ground when high-order ToM is required.
-2 Yōkai
+    We demonstrate that algorithm rankings change between HLE and YLE, suggesting that apparent progress in ZSC may be benchmark-specific.
+2 Background: Yōkai
 Figure 2: A sample round in nine-card YLE.
-Agent Refer to caption first observes two cards (one blue, one green), moves the blue card, and finishes by either (4a) revealing or (4b) placing a hint.
+Refer to caption first observes two cards privately (one blue, one green), moves the blue card, and finishes by either (4a) revealing or (4b) placing a hint.
 Refer to caption places a hint.
-Then, Refer to caption can start its turn.
+Then, Refer to caption can start their turn.
 Refer to caption chooses to end the game.
 Ending the game will return the final reward based on the outcome.
-This game is lost because while blue cards are clustered, green and red ones are not.
+This game is lost because green and red cards are not clustered.
+Figure 3: Legal moves for the card with the blue dot. Cards can only be moved so that all cards remain connected via their sides. Cards with an orange dot cannot currently be moved.
 Yōkai [^1] Detailed rules for Yōkai are available at http://boardgame.bg/yokai%20rules.pdf.
-is a turn-based cooperative game in which players must sort face-down cards into same-colour clusters to win.
+developed by Julien Griffon, is a turn-based cooperative game in which players must sort face-down cards into same-colour clusters to win.
 A key requirement is that all cards must be connected to at least one other card on their sides.
 Classically, the game is played by two to four players with 16 cards of four colours each.
-Complementing these 16 Yōkai cards, additional hint cards feature one or multiple colours and are used to hint at the colour of a face-down card underneath them.
-They are the only allowed communication mechanism in the game.
-The number of players determines the number of hint cards, which varies between seven and 10 (see Appendix B for further details).
-Hint cards are initially face-down and must be revealed before being placed face-up on a Yōkai card.
-Importantly, this Yōkai card is locked in place and can no longer be moved or observed.
+Complementing these Yōkai cards, additional hint cards feature one or multiple colours and are used to hint at the colour of a face-down card underneath them.
+Hint cards are the only allowed communication mechanism in the game, and their number depends on the number of players (see supplementary material; typically 4 – 10).
+These cards are initially face-down and must be revealed before being placed face-up on a Yōkai card.
+Importantly, a hinted Yōkai card can no longer be moved or observed.
 Since most hints can be multi-coloured, their interpretation is left to the other agents, depending on the context and history.
@@
 The current player first observes two cards privately, then moves one card, and then either (a) reveals or (b) places a face-up hint card.
-Consequently, the available legal moves depend on the specific card to be moved.
-Figure 11, while Figure 10 show legal moves and card configurations.
+The available legal moves depend on the specific card to be moved (an example is shown in Figure 3).
 The game ends when all hint cards have been revealed and placed.
-Optionally, players can end the game prematurely instead of observing cards at the beginning of their turn.
+Optionally, players can end the game early instead of observing cards at the beginning of their turn.
 The final game score is based on the final card configuration:
-$\begin{split}S=&\;5N_{\text{hints\_face\_down}}+2N_{\text{hints\_not\_placed}}+N_{\text{hints\_correct}}\\ &-N_{\text{hints\_wrong}}.\end{split}$ (1)
-Players are encouraged to end the game early to achieve a higher score, since unused hint cards contribute most to the reward.
-However, the earlier a player ends a game, the less information they can acquire first-hand and thus must infer knowledge from others’ behaviour (see Figure 12 for an example).
-As such, the final game score measures players’ joint ToM reasoning abilities.
+$\begin{split}S=5N_{\text{hints\_face\_down}}+2N_{\text{hints\_not\_placed}}+N_{\text{hints\_correct}}-N_{\text{hints\_wrong}}.\end{split}$ (1)
+Players are encouraged to end the game early to achieve a higher score, since unused hint cards contribute most to the final score.
+However, the earlier a player ends a game, the less information they can acquire first-hand and instead from others’ behaviour (see Figure 4 for an example).
+As such, the final game score reflects players’ joint ToM reasoning abilities.
 In particular, the rate at which players successfully end the game early reflects how much they rely on belief inference, making it a powerful new implicit measure of ToM reasoning under uncertainty.
+Figure 4: A second-order ToM reasoning example over multiple timesteps in 2-player YLE.
+Refer to caption beliefs that Refer to caption beliefs that they (Refer to caption ) knew that cards 1 and 2 are of the same colour.
+Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground.
+In the future, they now can potentially finish early as they have one less card that needs to be observed.
 Yōkai poses three key challenges to cooperative agents: First, achieving a high score requires agents to end a game as early as possible, resulting in a higher risk of failure.
-Second, ending the game early requires interpreting others’ actions and tracking their beliefs to form correct own beliefs over the colours of unobserved cards (see Figure 1).
+Second, ending the game early requires interpreting others’ actions and tracking their beliefs to form correct own beliefs over the colours of unobserved cards.
 Third, agents must use hints efficiently as grounded communication, but a single misplaced hint can render the game unwinnable.
 3 Preliminaries
-We model Yōkai as a decentralised partially observable Markov decision process (Dec-POMDP) Oliehoek and Amato (2016), formally described as a tuple $\mathcal{M}=\langle\mathcal{N},\mathcal{S},\Omega,O,A,\mathcal{R},\mathcal{T},\gamma\rangle$.
-$\mathcal{N}$ is the set of agents where $i\in\mathcal{N}$, $1\leq i\leq|\mathcal{N}|$ denotes one specific agent and i-i denotes all other agents.
-Set $\mathcal{S}$ denotes the possible true states of the environment, with $s\in\mathcal{S}$ being a single state.
-At timestep $t$ an agent in state $s_{t}$ receives a partial observation $o^{i}_{t}\in\Omega$ through the observation function $O(s_{t},i)$.
-The action-observation history per agent is denoted by $\tau^{i}$.
-Following Sclar et al.
-Sclar et al.
-(2022), an agent with no ToM chooses an action $a^{i}\in A$ only based on their observation of the environment $\pi^{i}(a^{i}|o^{i}_{t})$.
-Agents with zeroth-order ToM incorporate memory, conditioning their policy on a hidden state $h^{i}_{t}$ in addition to their current observation $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t})$.
-Furthermore, agents that reason about the mental states of others estimate their hidden states $\hat{h}_{t}^{(j)}$ for $j\neq i$ and condition their action on all $\hat{h}_{t}^{(-i)}$: $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t},\hat{h}_{t}^{-i})$.
-Based on the joint action performed by all agents $\bm{a}=(a_{1},\dots,a_{n})$ the environment returns a shared reward $r_{t}=R(s_{t},\bm{a})$ and then produces a new state based on the state transition function $\mathcal{T}$.
-Finally, using the discount factor $\gamma^{t}$, the agents optimise the discounted return of a trajectory given by $R(\tau)=\sum_{t}\gamma^{t}r_{t}$.
+Definition 3.1 (Dec-POMDP (Oliehoek & Amato, 2016)).
+[…]
+$J_{\text{SP}}(\pi)=\mathbb{E}_{\tau_{T}\sim\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\mathcal{R}(s_{t},\bm{a}_{t})\right].$ (2)
+Zero-Shot Coordination (ZSC) (Hu et al., 2020) aims to design algorithms that produce compatible policies in Dec-POMDPs, in that independently trained agents can coordinate.
+The compatibility between different joint policies $\pi_{1},...,\pi_{n}$ in an $n$-player Dec-POMDP is measured by the cross-play (XP) return, defined as in Forkel et al.
+(2026) as
+$J_{\text{XP}}(\pi_{1},...,\pi_{n}):=\mathbb{E}_{\phi\sim\text{Perm}(|\mathcal{N}|)}\left[J_{\text{SP}}((\pi_{\phi(1)}^{1},...,\pi_{\phi(n)}^{n}))\right],$ (3)
+where $\mathrm{Perm}(n)$ denotes the set of permutations of $\{1,\dots,n\}$.
+A common failure mode in ZSC is that agents learn partner-specific conventions that do not generalise to new partners, which results in XP returns being significantly lower than SP returns.
+Approaches to mitigate this SP-XP gap include Other-Play (OP) (Hu et al., 2020) and Off-Belief Learning (OBL) (Hu et al., 2021).
+Here, we focus on the commonly used inter-seed XP setting, where the implementation is fixed but random seeds differ (Hu et al., 2020; 2021; Cui et al., 2022; Muglich et al., 2025; Forkel et al., 2026).
 4 The Yōkai Learning Environment
-The Yōkai Learning Environment implements the game logic with two exceptions: First, we support configurable game sizes (e.g., a 3x3 version with nine cards and three colours).
-We refer to it as 3x3 while the original game is referred to as 4x4.
-The 3x3 two-player game has four hint cards: one one-colour hint and three two-colour hints.
-Second, our game is restricted to a $g\times g$ grid to efficiently compute legal moves.
-We pick $g$ large enough not to limit gameplay in practice and confirm this experimentally, see Appendix B.
-Our implementation is based on JaxMARL Rutherford et al.
-(2023) and is fully GPU-accelerated.
-It allows end-to-end learning on a single GPU at hundreds of thousands of steps-per-second, see Table 3.
+The Yōkai Learning Environment (YLE) implements a parametrised version of the collaborative card game Yōkai.
+We support configurable game sizes (e.g., 9C with nine cards and three colours, and 16C, the original game) and restrict play to a $g\times g$ grid to enable efficient computation of legal moves.
+The grid size $g$ is chosen large enough not to constrain gameplay in practice (see supplementary).
+The environment is implemented in JAX in JaxMARL (Rutherford et al., 2023) and supports end-to-end GPU training at hundreds of thousands of steps per second (see Table 1).
 Yōkai as a Graph
-We opted to model Yōkai as a graph since it is a natural formulation of the relational nature of the environment and allows for an efficient implementation in JAX.
-At time $t$, the state is a graph $G_{t}(V,A)$ with cards and hints as vertices $V=Y\cup H$ and adjacency matrix $A$.
-Within $G_{t}$, cards are vertices $V=Y\cup H$ and connected to neighbours via edges represented by an adjacency matrix $A$.
-Cards and hints are identified by their index $0\leq i<|Y|$ or $0\leq j<|H|$.
-Hint cards are not connected to any card.
-For example, in the 3x3 version with nine cards and four hints, colours (e.g., blue, green, red) and hints are randomly assigned each episode, and the graph $A\in\{0,1\}^{13\times 13}$ is updated when cards move.
-With this setup, moving a card $i$ boils down to obtaining $A^{\prime}$ from $A$ via $A^{\prime}=\text{move}(A,i,n)$ where in move the row and column vector at index $i$ of $A$ are replaced with a binary encoding of the new neighbours $n$.
-Additionally, a configuration is legal if all cards are connected, computed via a reachability matrix over $A$.
-Determining whether the game has been won then boils down to computing
-$\bigwedge\limits_{c\in C}\texttt{isConnected}(A_{c})$ (2)
-where $A_{c}$ denotes the submatrix of $A$ obtained by deleting all column and row vectors not associated with colour $c$, i.e.
-by finding the set of indices $\{j|0\leq j<|Y|\land Y[j]\neq c\}$ where $Y[j]$ denotes accessing the tuple $Y$ at index $j$.
-To compute the state-dependent set of legal move actions $A(s)\subseteq A$, we run $\texttt{isConnected}(A^{\prime})$ on all $A^{\prime}$ obtainable by moving a card to a position in the $g\times g$ field.
+We model Yōkai as a graph to reflect its relational structure and enable a efficient JAX implementation.
+At time $t$, the state is a graph $G_{t}(V,A)$ where vertices $V=Y\cup H$ represent cards and hints, and $A$ is the adjacency matrix encoding spatial neighbourhoods between cards.
+Hint cards are not connected to any vertex.
+Card colours and hints are randomly assigned at the beginning of each episode.
+Moving a card updates the corresponding row and column of $A$ to reflect its new spatial neighbours.
+A configuration is legal if all cards remain connected, which we verify via reachability on $A$.
+The game is won if, for each colour $c\in C$, all cards of colour $c$ form a connected component.
+Legal move actions are computed by evaluating connectivity after candidate moves.
 Observation Space
-Agents receive graph-based or image-like observations encoding card properties (e.g., colour, position, lock state) and hint features.
-Observations include both public (topology and hints) and private (card features) information, leading to asymmetric knowledge.
-Graph observations encode cards as an adjacency matrix $A_{t}\in\{0,1\}^{|Y|\times|Y|}$, and their properties as node features $F_{t}\in\mathbb{R}^{|Y|+|H|\times f}$, where $f$ is the feature dimension.
-In image-like observations, card features $F$ are embedded into a $g\times(g+1)\times f$ tensor based on their positions, while hint card features are embedded into column $g+1$.
-Agents observe which cards others inspect, but not their content.
-Partial observability forces agents to seek information and track shared knowledge.
+Agents receive either graph-based or image-like observations encoding card properties (e.g., colour, position, lock state, id) and hint features.
+Observations contain both public information (topology and hints) and private information (card features), leading to asymmetric knowledge.
+In the graph representation, cards are encoded by an adjacency matrix $A_{t}\in\{0,1\}^{|Y|\times|Y|}$ and node features $F_{t}\in\mathbb{R}^{(|Y|+|H|)\times f}$.
+In the image representation, card features are embedded into a $g\times(g+1)\times f$ image-like tensor according to spatial position.
+Agents observe which cards others inspect, but not their content, requiring belief tracking over time.
 Action Space
-Actions in the YLE are represented by a categorical action space.
-Additional players increase the number of hints and, thereby, the number of possible actions.
-The action space includes observing pairs of cards, moving cards, placing or revealing hints, ending the game, and a no-op.
-The action space size is the sum of the number of cards to look at $|Y|$, the number of target positions any card can be moved to $N_{moves}=g^{2}|Y|$, and the number of hints $|H|$.
-The number of cards a hint can be placed onto $N_{hint\_moves}=|H|\cdot|Y|$, $1$ action for no operation and $1$ action to end the game.
-The exact size of the action space depends on the setup but ranges from $\sim 1,000$ to $\sim 3,000$ in practice, all details are in Appendix B.
+Actions are represented categorically and include observing cards, moving cards, placing or revealing hints, ending the game, and a no-op.
+The action space size depends on the configuration and scales with the number of cards $|Y|$, grid positions $g^{2}$, and hints $|H|$.
+In practice, it ranges from approximately 1,0001{,}000 to 3,0003{,}000 actions (see supplementary).
+Invalid actions are masked at each timestep.
+Even with masking, this action space is substantially larger than those of widely used cooperative MARL and ZSC benchmarks (Bard et al., 2020; Carroll et al., 2019; Gessler et al., 2025) and therefore challenges existing algorithms with a larger exploration problem.
 Reward
-The score in Equation 1 is only given when the game ends successfully, and sometimes can be negative due to hint misplacements.
-Agents, however, may learn that it is better to avoid grouping cards altogether to avoid penalties.
-To address this, we modify the reward function.
-If the game ends without a win, we give a small negative reward to discourage this strategy:
-$R=\begin{cases}S&\text{if game won}\\ -d_{a}-(|C|-c_{a})-i_{h}&\text{otherwise.}\end{cases}$ (3)
-where $c_{a}$ is the number of correctly formed clusters, $i_{h}$ is the number of incorrectly placed hints, and $d_{a}\in\{0,1\}$ indicates whether the game was ended early.
-Since rewards are sparse, we shaped only the training reward to accelerate learning.
-We additionally reward observing a previously unobserved card (+1+1), grouping a single colour, and placing a correct hint (both +1+1) or an incorrect hint (1-1).
-5 Baselines, Experiments, and Results
-To evaluate the spatio-temporal belief tracking capabilities of agents in the YLE, we explore five complementary research questions (RQs).
-We begin by comparing agents with different neural architectures to establish baselines, then progressively test their capacity for belief modelling, generalisation, and reasoning in increasingly complex scenarios:
-RQ1
-How well do different policies reason over their memory (zero-order ToM)?
-RQ2
-Does learning belief representations and explicit action encoding improve agents’ performance?
-RQ3
-Do agents learn belief‑tracking strategies that let them coordinate with new partners?
-RQ4
-Do agents represent beliefs in their internal activations?
-RQ5
-Can agents scale to higher-order ToM scenarios involving four players?
-Setup
-We use the two-player 3x3 YLE version, which already presents a significant challenge for RL agents.
-Agents are trained using MAPPO Yu et al.
-(2022) for $10^{9}$ steps and evaluated on 1,000 games, across three random seeds.
-The maximum episode length was $32$ for this setting.
-We report four metrics: return (R), success rate (SR), number of colours clustered (NC) and the rate of successfully ending early (SEE).
-SEE captures how often agents end the game early and still win – a decision that requires strong belief inference under uncertainty that comes with significant risk of failure.
-We tuned key hyperparameters for our models (e.g., encoder size, depth, etc.); Training details are in Appendix C.
-5.1 RQ1a: Perfect memory is not sufficient to solve the YLE
-Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
-Human - $5.7\pm 0.2$.5) $95.0\pm 2.5$.8) $3.0\pm 0.0$.0) $65.0\pm 4.8$.8)
-GCN-M perfect $0.9\pm 0.0$.5) $26.8\pm 1.3$.4) $1.6\pm 0.0$.3) $0.0\pm 0.0$.0)
-GATv2-M perfect $1.9\pm 0.0$.4) $58.2\pm 1.3$.5) $2.3\pm 0.0$.2) $0.0\pm 0.0$.0)
-RM-M perfect $1.6\pm 0.1$.1) $50.7\pm 3.5$.0) $2.2\pm 0.0$.6) $0.0\pm 0.0$.0)
-CNN-M perfect $2.3\pm 0.0$.3) $71.8\pm 0.8$.8) $2.6\pm 0.0$.1) $0.0\pm 0.0$.0)
-CNN-LSTM implicit $0.2\pm 0.0$.1) $3.9\pm 0.4$.1) $0.8\pm 0.0$.3) $1.0\pm 0.1$.8)
-CNN-GRU implicit $0.3\pm 0.0$.2) $2.3\pm 0.0$.3) $0.7\pm 0.0$.1) $2.1\pm 0.1$.8)
-CNN-S5 implicit $0.1\pm 0.0$.1) $3.0\pm 0.3$.8) $0.7\pm 0.0$.2) $0.1\pm 0.0$.2)
-CNN-TrXL explicit $1.5\pm 0.0$.3) $43.5\pm 0.9$.7) $2.1\pm 0.0$.2) $0.0\pm 0.0$.0)
-ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
-Table 1: Agent performance on the 3x3 YLE.
-The first block compares encoder architectures with access to perfect memory (-M).
-The second block evaluates different temporal architectures with the best-performing CNN encoder.
-There, we compare implicit and explicit memory modules.
+The game returns the final score only at termination; the score can be negative due to incorrect hint placements.
+To discourage trivial strategies (e.g., never attempting to win to avoid hint penalties), we modify the terminal reward from Equation 1:
+$R=\begin{cases}S&\text{if the game is won},\\ -d_{a}-(|C|-c_{a})-i_{h}&\text{otherwise},\end{cases}$ (4)
+where $|C|$ is the number of colours, $c_{a}$ is the number of correct clusters, $i_{h}$ the number of incorrect hints (colour not included in the hint), and $d_{a}\in\{0,1\}$ indicates early termination.
+We additionally apply reward shaping to accelerate learning (rewarding newly observed cards and correct hints, both +1+1 each), which is annealed to zero during training and disabled at evaluation time.
+Successful Early Ending (SEE)
+Table 1: We evaluate steps-per-second (SPS) of YLE on a Nvidia H100 by taking random actions in $n$ parallel environments. YLE scales to hundreds of thousands of SPS.
+Num Envs $n$ 512 1,024 2,048
+2-player 9C 139,107 274,669 559,251
+2-player 16C 135,662 144,636 148,269
+Because unused hint cards yield the largest reward, every additional hint implicitly reduces the maximum attainable score, creating strong time pressure to end the game once sufficient common ground has been established.
+We therefore introduce Successful Early Ending (SEE) as an additional metric next to reward.
+SEE is the rate at which agents terminate the game early and still win.
+It is complementary to average game length: long games indicate overly cautious play, while short but unsuccessful games reflect guessing without sufficient belief formation.
+This positions SEE as a diagnostic signal of functional ToM reasoning.
+Implementation
+Our Jax-based implementation scales efficiently on GPU, reaching hundreds of thousands of steps per second (see Table 1) and performing comparably to other JAX-based RL environments (Rutherford et al., 2023; Nikulin et al., 2023; Matthews et al., 2024; Ruhdorfer et al., 2025b).
+5 Adapting ZSC Methods to YLE
+5.1 Other-Play in YLE
+Other-Play (Hu et al., 2020) modifies self-play by optimizing expected return under random symmetry transformations drawn from the known symmetries of the Dec-POMDP.
+In practice, OP is implemented as domain randomization: at the start of each episode, each agent samples a transformation and interacts with a permuted view of the same underlying environment.
+This discourages convergence to arbitrary symmetry-breaking conventions that fail under cross-play.
+In Hanabi, the dominant symmetry corresponds to permutations of card colours.
+Since card positions carry no semantic meaning and there is no spatial structure, enforcing invariance to colour permutations is sufficient to prevent most seed-dependent conventions.
+In the YLE, symmetry structure is richer due to the spatial layout and card movement.
+Symmetry breaking can arise from both card colours and the spatial configuration.
+Enumerating all Dec-POMDP symmetries in complex environments is generally intractable in practice, which is a known limitation of OP.
+We therefore focus on the two most salient symmetry families that directly induce symmetry-breaking: (i) colour permutations and (ii) spatial grid rotations.
+Both admit symmetry breaking, e.g., clustering specific colours in fixed regions.
+We thus define the symmetry group as $\Phi=\Phi_{\text{col}}\times\Phi_{\text{rot}}$, where $\Phi_{\text{col}}$ permutes colours and $\Phi_{\text{rot}}=\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ rotates the grid.
+Unlike Hanabi, multiple symmetry classes exist in the YLE. Using only a subset may therefore leave residual symmetry-breaking conventions, which we show empirically in section 6.
+5.2 Off-Belief Learning in YLE
+Off-Belief Learning (Hu et al., 2021) constructs a sequence of policies and belief models.
+At level $\ell$, policy $\pi_{\ell}$ chooses actions that are optimal assuming (i) past actions were generated by a reference policy $\pi_{\ell-1}$, used to form beliefs, and (ii) future actions will follow $\pi_{\ell}$.
+Iterating this operator yields a hierarchy $\pi_{0}\rightarrow\pi_{1}\rightarrow\dots\rightarrow\pi_{L}$ that controls reasoning depth and, produces a unique policy per level given $\pi_{0}$.
+OBL relies on a learned belief model $\mathcal{B}_{\pi}$ that predicts the unknown state-action history $\tau_{t}$ given a joint policy $\pi$ and an agent’s action-observation history $\tau_{t}^{i}$, i.e.
+$\mathcal{B}_{\pi}(\tau_{t}|\tau_{t}^{i})$.
+In Hanabi, this belief predicts hidden card identities in the observing player’s hand.
+In YLE, the (private, unknown) state consists of the unobserved colours of face-down Yokai cards.
+Accordingly, we train $\mathcal{B}_{\pi}$ to predict a posterior over the global board configuration conditioned on the agent’s AOH.
+In doing so, we closely follow the methodology of Hu et al.
+(2021): Our belief model features an LSTM (Hochreiter & Schmidhuber, 1997) encoder and autoregressive decoder that predicts a colour per card slot (i.e.
+nine slots for nine cards).
+Cards have a unique id associated with them in the observation which makes it possible for the belief model to map cards in observations to these slots.
+Per slot, we ensure that cards are sampled consistently with the ego agents AOH: We respect the global card colour restrictions (i.e.
+9 cards of 3 colours) and only sample cards which are unknown first-hand to the agent from $\mathcal{B}$.
+To our knowledge this is the first implementation of OBL outside of Hanabi.
+Implementing OBL outside of Hanabi highlights some limitations of OBL as formulated in Hu et al.
+(2021).
+First, hints in the YLE are not guaranteed truthful.
+The belief model must therefore treat hints as unreliable evidence, making posterior estimation more challenging than in Hanabi.
+Second, OBL theory requires a fictitious rollout length $K$ such that the acting agent moves again.
+While two-player Hanabi requires $K=2$, YLE’s multi-step turn structure requires $K=4|\mathcal{N}|$, increasing fictitious rollout cost substantially.
+In practice, this makes OBL expensive relative to other ZSC methods both in terms of additional environment steps and memory requirements.
+Such limitations are highlighted by testing ZSC algorithms in additional environments, a contribution of this work.
+6 Experiments
+We evaluate the YLE as a benchmark for zero-shot coordination by comparing self-play (SP) and cross-play (XP) performance across leading ZSC methods.
+We first show that the YLE is solvable in self-play, with agents achieving super-human performance in the simplest settings.
+However, strong self-play does not translate into strong cross-play.
+Across methods that achieve near-perfect inter-seed XP in Hanabi, we observe persistent SP–XP gaps and substantial degradation in XP performance in the YLE, indicating that existing ZSC approaches do not generalise to its demands.
+6.1 Setup
+Agents are trained using independent PPO (IPPO) (de Witt et al., 2020) for $10^{9}$ steps and evaluated on 5,000 games.
+Our default policy consists of a 4-layer CNN encoder (64 filters, kernel size $(3,3)$, ReLU, stride 1), followed by a linear projection, a GRU, and actor–critic heads.
+All hidden dimensions are 256.
+For OBL, we adopt the standard public–private architecture (Hu et al., 2021; Cui et al., 2022) with separate CNN encoders for private and public observations.
+We optimise with Adam (Kingma & Ba, 2015) ($\epsilon{=}10^{-5}$), $\gamma{=}0.99$, $\lambda{=}0.95$, value loss coefficient $0.5$, four PPO epochs with four mini batches of 128 steps, and linearly annealed reward shaping.
+We tune the PPO entropy coefficient $\alpha$ per setting, sweeping $\alpha\in[0.01,0.08]$.
+We report results over 12 random seeds per configuration.
+We report return (R), successful early ending (SEE), and game length (LEN).
+SEE measures how often agents terminate early and still win, and is decomposed into early-ending rate (EE) and calibration $\textbf{WEE}=P(\text{W}|\text{EE})$.
+Our primary comparison metric is cross-play return.
+Maximum scores and episode lengths vary with the number of players and environment size (9C vs 16C) due to the differing number of available hints22244, 55, 66 for 2, 3, and 4 players respectively in 9C; $7$, 99, $10$ in 16C..
+In 9C YLE, the maximum episode length is $32$, $40$, and $48$, and the theoretical maximum score is $20$, $25$, and $30$ for 2, 3, and 4 players respectively.
+Achieving this maximum requires immediate early termination without interacting with the board.
+By contrast, a policy that always uses all hints correctly and never terminates early obtains only $4$, 55, and $6$ respectively.
+Further details are provided in subsection C.2 in the supplementary.
+6.2 Zero-Shot Coordination Experiments
+Table 2: CNN-GRU SP and XP performance under different algorithms and settings in the YLE with humans (H).
+To isolate coordination challenges from memory limitations we use a perfect memory setting (M; indicated by ) where we show agents the values of already seen cards even if they have been moved.
+Full YLE is indicated by .
+Each entry in table is the mean $\pm$ standard error of the mean computed from $12$ independent seeds.
+The table shows $180$ trained policies.
 Best results are in bold.
-ToMTrack (final block) is the best method overall when not given perfect memory.
-Figure 3: Training curves for the 3x3 YLE, with bands showing the standard error over three seeds. Left: comparison of encoders with perfect memory. Right: comparison of memory architectures.
-To evaluate if agents can reason about their own memory (zero-order ToM), we first compared different encoder architectures in a perfect memory setting: In this setting, agents were given perfect memory by including all previously observed cards in their observations.
-This experiment also served to set a performance upper bound.
-Given the YLEs spatial reasoning task and graph-based nature, we compared CNN-, GCN- Kipf and Welling (2016), GATv2- Brody et al.
-(2021), and Relation Module (RM)-based encoders Zambaldi et al.
-(2018).
-Architectural details are provided in Appendix C.2.
+The column headers are: V = Card version (9C, 16C) and P = # of players (2, 3, 4).
+The difference between highest average XP and SP return is 7.54.8=2.77.5-4.8=2.7 (a $43.9\%$ difference).
+Self-Play (SP) Cross-Play (XP)
+Alg. V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
+H 9C 2 $5.7\pm 0.0$.5) $65\pm 11.$ $65\pm 11.$ $100\pm 0.$
+SP 9C 2 $7.5\pm 0.0$.3) $89\pm 7.$ $89\pm 8.$ $92\pm 8.$ $25.0\pm 0.0$.6) $1.6\pm 0.0$.2) $18\pm 3.$ $35\pm 5.$ $50\pm 3.$ $29.7\pm 0.0$.4)
+HE 9C 2 $7.5\pm 0.0$.1) $95\pm 0.$ $6\pm 0. 99$$9\pm 0$. $24.6\pm 0.0$.1) $4.1\pm 0.0$.8) $50\pm 1.$ $68\pm 10.$ $69\pm 7.$ $27.2\pm 0.0$.7)
+OP 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
+OBL1 9C 2 $4.0\pm 0.0$.2) $68\pm 7.$ $76\pm 8.$ $83\pm 7.$ $28.0\pm 0.0$.4) $2.3\pm 0.0$.3) $43\pm 7.$ $56\pm 9.$ $76\pm 2.$ $29.4\pm 0.0$.4)
+OBL2 9C 2 $4.3\pm 0.0$.2) $69\pm 7.$ $79\pm 7.$ $81\pm 7.$ $27.7\pm 0.0$.4) $2.1\pm 0.0$.2) $38\pm 4.$ $50\pm 5.$ $75\pm 2.$ $29.5\pm 0.0$.2)
+OBL3 9C 2 $4.4\pm 0.0$.4) $73\pm 7.$ $82\pm 7.$ $81\pm 8.$ $27.3\pm 0.0$.4) $2.2\pm 0.0$.3) $41\pm 6.$ $55\pm 6.$ $73\pm 3.$ $29.1\pm 0.0$.4)
+OBL4 9C 2 $4.5\pm 0.0$.2) $77\pm 3.$ $87\pm 3.$ $89\pm 2.$ $27.2\pm 0.0$.2) $2.8\pm 0.0$.3) $53\pm 8.$ $66\pm 8.$ $78\pm 3.$ $28.5\pm 0.0$.5)
+OBL5 9C 2 $4.8\pm 0.0$.2) $80\pm 3.$ $89\pm 2.$ $90\pm 2.$ $27.0\pm 0.0$.2) $2.5\pm 0.0$.3) $47\pm 7.$ $61\pm 7.$ $75\pm 3.$ $28.9\pm 0.0$.3)
+SP 9C 3 $7.8\pm 0.0$.6) $64\pm 10.$ $67\pm 11.$ $73\pm 12.$ $33.4\pm 0.1$.1) $0.2\pm 0.0$.1) $0\pm 0.$ $6\pm 1.$ $17\pm 1.$ $39.5\pm 0.0$.1)
+HE 9C 3 $0.9\pm 0.0$.0) $3\pm 0.$ $100\pm 0.$ $3\pm 0.$ $1.0\pm 0.0$.0) $0.9\pm 0.0$.0) $4\pm 0.$ $100\pm 0.$ $4\pm 0.$ $1.0\pm 0.0$.0)
+OP 9C 3 $3.0\pm 0.0$.8) $27\pm 0.$ $38\pm 11.$ $37\pm 11.$ $36.2\pm 0.1$.1) $1.2\pm 0.0$.2) $10\pm 1.$ $29\pm 4.$ $34\pm 3.$ $38.0\pm 0.0$.2)
+SP 9C 4 $0.5\pm 0.0$.2) $3\pm 1.$ $50\pm 15.$ $3\pm 1.$ $38.1\pm 0.3$.0) $0.1\pm 0.0$.0) $1\pm 0.$ $64\pm 20.$ $1\pm 0.$ $36.6\pm 0.3$.7)
+HE 9C 4 $0.7\pm 0.0$.1) $3\pm 0.$ $100\pm 0.$ $3\pm 0.$ $17.0\pm 0.3$.9) $0.7\pm 0.0$.0) $2\pm 0.$ $100\pm 0.$ $2\pm 0.$ $6.6\pm 0.2$.1)
+OP 9C 4 $0.2\pm 0.0$.0) $1\pm 0.$ $49\pm 14.$ $1\pm 0.$ $38.8\pm 0.2$.6) $0.2\pm 0.0$ $1\pm 0.$ $68\pm 6.$ $2\pm 0.$ $36.5\pm 0.1$.5)
+SP 16C 2 $0.0\pm 0.0$.0) $0\pm 0.$ $25\pm 12.$ $0\pm 0.$ $52.6\pm 0.1$.8) $0.0\pm 0.0$.0) $0\pm 0.$ $25\pm 16.$ $0\pm 0.$ $53.2\pm 0.1$.7)
+As introduced, we compare SP, high entropy IPPO (HE) , OP, and OBL in a perfect-memory setting to isolate coordination challenges.
+For HE, we follow Forkel et al.: in 2P we use $\lambda_{\text{GAE}}=0.85$ and $\alpha=0.07$, and in multi-player settings we sweep $\alpha\in\{0.01,0.02,\dots,0.07\}$.
+For other 2P settings we use $\alpha=0.05$, while for 3P and 4P we use $\alpha=0.01$.
+We verify these choices via ablations below.
+Our final results are the result of training and comparing $400$ policies.
 To anchor results, we additionally establish a human performance baseline based on results from $25$ games played by $5$ participants (ages 24 - 28; see Appendix F).
-As shown in Table 1 (top), all agents severely underperform the human baseline.
-Among agents, despite YLEs graph-based nature, the CNN encoder outperformed both GNNs and the RM in all metrics except SEE (see Figure 3, left).
-We hypothesise this is due to YLE’s dynamic spatial layout: card positions and relationships change frequently, which may be better captured by CNNs than by message-passing.
-In contrast to humans (SEE 65%), agents failed to solve the task reliably or finish early.
-This suggests that the YLE requires more than memory: It demands the ability to reason about asymmetric knowledge to maintain common ground over time and to act under uncertainty.
-5.2 RQ1b: Explicit memory modules show stronger zero-order ToM than implicit ones
-While the perfect memory setting establishes an upper bound, it assumes agents retain all past observations without limitation.
-To better reflect realistic constraints and to asses whether agents can reason over their own internal knowledge representations, we evaluated models with different temporal architectures.
-This tests their capacity to track and use information over time – a key aspect of zero-order ToM.
-We tested LSTM- Hochreiter and Schmidhuber (1997), GRU Cho et al.
-(2014), S5- Smith et al.
-(2023), and TransformerXL-based (Dai et al., 2019, TrXL) architectures, paired with the best-performing CNN encoder.
-S5 has recently excelled on memory tasks in Samsami et al.
-(2024), while TrXL includes an explicit memory mechanism shown effective in memory-intensive RL Parisotto et al.
-(2020).
-As shown in Table 1 (middle) and Figure 3 (right), TrXL outperformed all other temporal models ($p<0.05$, one-sided paired t-test).
-Nonetheless, there is still a performance gap between TrXL and the best perfect memory baseline CNN-M.
-Contrary to expectations from prior work Samsami et al.
-(2024), S5 offered no clear advantage over traditional RNNs.
-LSTM, GRU, and S5 agents occasionally ended games early.
-This behaviour appears to reflect a form of gambling, as agents terminate the game after very few actions and without sufficient information-gathering.
-This suggests a failure to track beliefs over time, as agents that end early do so without sufficient inference – contrasting with human players, whose early-ending behaviour reflects confident ToM reasoning.
-This suggests a failure to reliably recall past observations, leaving guessing as the primary strategy.
-This tendency is further reflected in the short average game length ($22.9\pm 15.5$, $14.0\pm 15.6$ and $30.4\pm 1.42$, respectively).
-TrXL agents never finish early.
-Taken together, these results indicate that explicit memory architectures, such as TrXL, are better suited for zero-order ToM reasoning.
-Nonetheless, the gap in performance relative to the perfect memory baseline demonstrates the difficulty of learning robust belief-tracking strategies in YLE.
-5.3 RQ2: Modelling beliefs over cards and distilling partner policies improves results
-Figure 4: Our ToMTrack agent generates beliefs over card colours using an auxiliary prediction head.
-Figure 5: Left: compares return. Right: shows success rate. ToMTrack trains at least $1.5\times$ faster.
-Given the partial observability of the YLE, agents must maintain a belief $b$ over the hidden colours $c$ of all cards.
-To support this, we designed ToMTrack, a method for learning an implicit model over cards’ colour (see Figure 4).
-ToMTrack is an online representation learning RL approach inspired by previous work on offline belief representation learning Wang et al.
-(2023).
-At each timestep $s_{t}$, the policy $\pi$ should learn useful representations that contain information about the cards’ colours, no matter if they have been previously observed or inferred (first-order ToM, recall Figure 1).
-We thus want our policy to model $p(c_{i}|s_{t}),\,\forall\,0<i<|Y|$.
-To achieve this, we added an auxiliary belief classification head $f_{\text{belief}}$ to the policy network.
-This head maps the intermediate representation learned by the policy to a distribution over possible card colours.
-During training, we used the ground truth card colours as labels.
-We used an auxiliary belief prediction task Jaderberg et al.
-(2017) to train $f_{\text{belief}}$.
-Specifically, we added a cross-entropy loss $\mathcal{L}^{\text{CE}}$ to the actor, $\mathcal{L^{\prime}}_{\text{a}}=\mathcal{L}_{\text{a}}+\lambda\mathcal{L}_{a}^{\text{CE}}$, and critic loss, $\mathcal{L^{\prime}}_{\text{c}}=\mathcal{L}_{\text{c}}+\lambda\mathcal{L}^{\text{CE}}_{c}$; where $\lambda\geq 0$ is a hyperparameter.
-In addition, since information is often implicitly communicated using actions, we add a second input encoder to explicitly process the last action taken by the active agent.
-This enables the policy to explicitly interpret observed behaviour and update its beliefs about card colours accordingly, see Figure 1.
-We also evaluated other baselines based on the simplified action decoder (SAD) Hu and Foerster (2020) or intrinsic motivation, but found that they underperformed, see Appendix H.
-In contrast, ToMTrack outperforms CNN-TrXL ($48.3$ vs $43.5$ SR) and decreases the gap to CNN-M ($71.8$ SR) (Table 1).
-These results demonstrate that modelling latent beliefs and interpreting others’ actions improves agent coordination.
-5.4 RQ3: Agents do not exhibit transferable models of partner belief inference
-Figure 6: Zero-shot coordination performance measured by SR.
-We report additional metrics in Appendix E.2.
-Self-play agents do not learn to coordinate with others.
-Other-Play agents (OP) have some success at partner generalisation.
-However, the additional complexity of YLE requires more symmetries as input to OP (+C+R) compared to Hanabi, which only requires +C.
-Figure 7: ToM reasoning influences Refer to caption ’s decision.
-Without ToM reasoning, no move is optimal.
-A zero-order agent (memory-based) may move either the red or blue card, as both would group two cards.
-A first-order agent that correctly interprets Refer to caption ’s action (see Figure 1) will recognise that moving the blue card completes a full cluster.
-Agents are rewarded during training for forming complete clusters.
-An example of second-order ToM reasoning can be found in Appendix B.3.
-Effective strategies for tracking one’s own and others’ beliefs should generalise to novel partners.
-We test this via zero-shot coordination (Hu et al., 2020, ZSC), where agents must cooperate with independently trained partners at test time.
-Success in ZSC indicates that agents do not rely on partner-specific conventions but instead learn to track beliefs, interpret actions, and reason about others’ knowledge in a generalisable way.
-We first evaluate ZSC performance of the previously trained CNN-M and ToMTrack agents using cross-play (XP), where agents are paired with separately trained partners.
-Figure 6 shows that XP SR drops by $88.5\%$ on average, indicating a strong reliance on learned conventions.
-To address this, we retrain agents using other-play (Hu et al., 2020, OP), which prevents overfitting to arbitrary conventions by exploiting environment symmetries.
-In Hanabi Bard et al.
-(2020), colour symmetry alone (OP+C) suffices for strong ZSC Hu et al.
-(2020).
-We find that this is not the case in YLE.
-As shown in Figure 6, while both CNN-M and ToMTrack trained with OP+C improve XP performance, a significant SP-XP performance gap remains.
-Unlike Hanabi, YLE introduces spatial-temporal complexity and thus additional symmetry classes (see Appendix C.4), e.g.
-agents could form conventions around locations.
-We therefore also test OP+C+R, using rotational symmetries to break spatial conventions.
-This improves XP further for CNN-M and ToMTrack.
-For CNN-M it nearly closes the SP-XP gap ($33.2$ vs $27.4$).
-However, while OP+C+R improves XP, this comes at a significant gap in overall SR (e.g.
-$86.5$ vs $41.5$ SP in CNN-M), suggesting that non-OP agents rely heavily on spatial and colour conventions.
-When these are unavailable, coordination suffers – implying that belief tracking and action interpretation remain difficult to learn directly.
-This suggests that CNN-M and ToMTrack do not acquire general belief strategies out-of-the-box, though OP-trained variants may begin to.
-Agent Memory T0 $\downarrow$ T1 $\downarrow$ Wrong $\uparrow$
-CNN-M-OP+C+R Perfect $13.8\pm 0.0$.5) $86.5\pm 1.4$.2) $30.3\pm 0.3$.1)
-ToMTrack-OP+C+R Explicit $12.6\pm 0.0$.9) $98.8\pm 2.8$.3) $28.2\pm 0.2$.7)
-Table 2: Results from the diagnostic ToM test in Figure 7.
-We report the average rank (based on action probabilities) of zero-order (T0), first-order (T1), and incorrect actions (e.g., mismatched card colours).
-We find that agents fail to make zero- or first-order inferences in this simple test.
-To further test agents’ ability to perform first-order ToM reasoning, i.
-e, to infer a partner’s beliefs and intentions, we designed a diagnostic task (see Figure 7).
-In this test, only agents with first-order ToM will correctly identify that they can group all blue cards with one move, despite not being able to observe all blue cards first-hand.
-Grouping blue cards is optimal as it increases the number of card clusters by one, increasing the chance of finishing successfully or possibly early.
-As shown in Table 2, we find that even the best-performing ZSC agents fail to select optimal actions and instead make moves consistent with either zero-order or no ToM reasoning.
-This shows that, despite OP improvements, agents lack transferable models of belief inference.
-In summary, agents in YLE struggle to track beliefs of novel partners.
-Unlike Hanabi, success in YLE likely cannot be achieved by symmetry-breaking alone: as additional symmetries are removed, performance declines.
-This suggests the bottleneck is not symmetry identification but rather the absence of learning mechanisms that foster generalisable ToM reasoning.
-Figure 8: Probing accuracies for each step. Memory Oracle observes new cards whenever possible and recalls them perfectly. Agents fail to track colours over time, especially after the mid-game.
-Figure 9: Performance comparison of the two-player (ToMTrack) and four-player (ToMTrack-4P) 3x3 YLE.
-4-player YLE requires tracking common ground between more players, and results in lower performance with higher variance.
-5.5 RQ4: Agents struggle to represent knowledge of self and others over longer timespans
-To succeed in the YLE, agents must maintain a belief over all cards’ colours.
-However, our results so far suggest that models are not using beliefs correctly.
-To identify whether agents fail to represent beliefs or instead only fail to use them, we apply linear probes Alain and Bengio (2016) to the agents’ hidden activations at each timestep of the game, to predict the colour of each card individually Matiisen et al.
-(2022); Bortoletto et al.
-(2024a, b); Zhu et al.
-(2024).
-This approach assesses whether belief-relevant information is linearly decodable from the learned embeddings – whether the card’s colour was directly observed, inferred, or remains uncertain (see Appendix D for details).
-Figure 9 shows that agents start to form internal representations accurately in the early-to-mid-game.
-However, performance degrades in the later stages.
-This suggests that agents struggle to retain and update beliefs over longer timespans, especially as cards move and the spatial arrangement grows more complex.
-Together, this indicates a challenge in representing dynamic beliefs in YLE.
-5.6 RQ5: Agents struggle to scale belief reasoning in four-player YLE
-To test whether agents can scale to higher-order Theory of Mind (ToM) reasoning, we evaluate them in the four-player version of YLE.
-In this setting, each player observes fewer cards directly, and knowledge is distributed across more teammates.
-As a result, coordination increasingly depends on inferring what others know and what others believe about each other’s knowledge.
-For example, reasoning about whether agent 1 understands what agent 2 knows about agent 3’s interpretation of agent 4’s hint involves third-order ToM.
-This demand is reflected in performance: as shown in Figure 9, ToMTrack exhibits increased variance in both SR and R, and mean SR drops by over $10$ percentage points compared to the two-player setting.
-Note that the four-player game allows higher absolute returns due to more hint cards (see Figure 9, left).
-However, as reflected in the success rate, this does not suggest better coordination.
-Our results suggest that scaling ToM reasoning in multi-agent RL remains an open challenge, especially in settings that require maintaining shared beliefs across multiple agents over time.
-5.7 Key takeaways & Limitations
-Across all RQs, agents, even with explicit memory or belief modelling, fail to maintain accurate beliefs or coordinate reliably over time.
-They overfit to partner-specific conventions and struggle with belief tracking over longer horizons.
-Notably, humans successfully end games early in 65% of cases, while agents very rarely do so – even with perfect memory or belief modelling (see Table 1).
-This finding positions SEE as a compelling metric for assessing belief reasoning, as it reflects agents’ ability to infer shared beliefs and identify coordination success.
-We present two main takeaways: First, ToM reasoning must be evaluated with novel partners to avoid mistaking brittle conventions for genuine ToM — ToM tests must acknowledge the generality constraint Rakoczy (2012).
-Second, YLE offers a challenging environment for assessing common ground reasoning under uncertainty, even in its simplest 2-player 3×3 setting.
-It supports future work to explore scaling to more agents and the original game size.
-Both enable an investigation of increasingly complex ToM settings.
-While the YLE significantly expands our ability to evaluate key elements of ToM and advance cooperative agents, it also has limitations: It currently abstracts away modalities used in human ToM, such as language or gaze, and focuses on isolating belief-based reasoning.
-Additionally, it focuses solely on cooperation and does not address belief modelling in competitive or mixed-motive settings.
-These simplifications allow for focused evaluation of core ToM challenges before addressing the full complexity of human ToM in more real-world settings.
+Our experiments run on NVIDIA V100, L40S, and H100 GPUs.
+A single seed takes between 6 and 12 hours depending on player and card count.
+Due to computational cost of training multiple levels and the methods performance, we evaluate OBL only in the 2P setting.
+We show XP matrices in Appendix H (Figures 15 - 24) and training curves in Appendix I which show our policies and belief models converging.
+Table 2 highlights four key results.
+First, state-of-the-art ZSC methods exhibit persistent SP–XP gaps, indicating failure to learn fully symmetric policies in the YLE.
+Despite the large discrete action space in YLE, agents consistently achieve strong self-play performance, indicating that the main difficulty arises from coordination.
+Second, XP performance remains below the human baseline despite agents having perfect memory.
+Third, HE — the state-of-the-art method for Hanabi – underperforms OP in the YLE and collapses in multi-player settings, often learning to terminate immediately ($\textbf{LEN}=1.0$ and $\textbf{LEN}=3.8$).
+Increasing entropy makes coordination impossible before reaching high XP scores, suggesting that entropy tuning alone is insufficient for YLE.
+OBL also underperforms OP.
+Notably, this reverses the ranking observed in Hanabi, where HE outperforms OP.
+Fourth, difficulty increases systematically with additional players and in the 16C setting, all of which remain unsolved even w.r.t.
+SP returns.
+These results establish the YLE as a challenging benchmark for ZSC and cooperative MARL.
+6.3 Ablations and Additional Experiments
+Table 3: We sweep OP PPO entropy coefficients $\alpha$ to verify our experimental setting.
+Self-Play (SP) Cross-Play (XP)
+$\alpha$ V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
+0.01 9C 2 $3.3\pm 0.0$.3) $32\pm 8.$ $41\pm 10.$ $46\pm 11.$ $29.3\pm 0.0$.6) $1.1\pm 0.0$.2) $12\pm 2.$ $28\pm 4.$ $47\pm 4.$ $30.2\pm 0.0$.4)
+0.03 9C 2 $6.2\pm 0.0$.1) $85\pm 1.$ $92\pm 1.$ $92\pm 1.$ $25.6\pm 0.0$.2) $4.2\pm 0.0$.1) $56\pm 2.$ $76\pm 3.$ $74\pm 1.$ $26.7\pm 0.0$.3)
+0.05 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
+0.08 9C 2 $0.1\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0) $0.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0)
+0.01 9C 3 $3.0\pm 0.0$.8) $27\pm 0.$ $38\pm 11.$ $37\pm 11.$ $36.2\pm 0.1$.1) $1.2\pm 0.0$.2) $10\pm 1.$ $29\pm 4.$ $34\pm 3.$ $38.0\pm 0.0$.2)
+0.03 9C 3 $1.0\pm 0.0$.1) $8\pm 4.$ $62\pm 13.$ $11\pm 6.$ $19.5\pm 0.5$.3) $0.1\pm 0.0$.1) $1\pm 0.$ $44\pm 14.$ $1\pm 0.$ $24.9\pm 0.5$.0)
+0.01 9C 4 $0.2\pm 0.0$.0) $1\pm 0.$ $49\pm 14.$ $1\pm 0.$ $38.8\pm 0.2$.6) $\pm 0.0 1±0.1\pm 0$. $68\pm 6.$ $2\pm 0.$ $36.5\pm 0.1$.5)
+0.03 9C 4 $0.1\pm 0.0$.1) $1\pm 0.$ $17\pm 11.$ $1\pm 0.$ $42.5\pm 0.3$.9) $0.1\pm 0.0$.1) $1\pm 0.$ $33\pm 10.$ $1\pm 0.$ $38.7\pm 0.3$.0)
+PPO Entropy Coefficient $\alpha$
+The PPO entropy coefficient $\alpha$ plays an important role in ZSC.
+Both because it improves SP performance due to the added exploration empirically in the YLE but also since high entropy coefficients improve XP performance.
+Specifically, Forkel et al.
+(2026) show that in Hanabi, near-perfect inter-seed XP can be achieved by picking a high enough $\alpha$ that maximises XP performance while not yet degrading the policy due to the added randomness.
+In Table 3, we sweep $\alpha$ for OP.
+In the 2P setting, moderate entropy improves XP performance, consistent with prior findings.
+However, higher entropy does not achieve state-of-the-art YLE inter-seed XP performance and is outperformed by OP.
+In 3P and 4P settings, increasing entropy rapidly destabilises learning, leading to policy collapse before coordination benefits can materialise.
+These results have two implications.
+First, even when combining symmetry-based training and entropy tuning, persistent SP–XP gaps remain in YLE.
+Second, unlike Hanabi, inter-seed XP in the YLE is not solvable via entropy threshold tuning alone.
+Other-Play Symmetries
+Table 4: OP with varying symmetry groups (G): full $\Phi$, only $\Phi_{\text{col}}$, and only $\Phi_{\text{rot}}$.
+Self-Play (SP) Cross-Play (XP)
+G V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
+$\Phi$ 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
+$\Phi_{\text{col}}$ 9C 2 $7.6\pm 0.0$.1) $97\pm 0.$ $97\pm 1.$ $100\pm 0.$ $24.7\pm 0.0$.1) $2.1\pm 0.0$.6) $24\pm 8.$ $39\pm 9.$ $55\pm 5.$ $29.5\pm 0.0$.7)
+$\Phi_{\text{rot}}$ 9C 2 $6.4\pm 0.0$.3) $77\pm 7.$ $82\pm 7.$ $86\pm 8.$ $27.0\pm 0.0$.5) $4.8\pm 0.0$.2) $63\pm 3.$ $74\pm 3.$ $85\pm 2.$ $27.0\pm 0.0$.3)
+As described in subsection 5.1, the YLE symmetry group factorises as $\Phi=\Phi_{\text{col}}\times\Phi_{\text{rot}}$.
+In Table 4, we compare OP trained with the full group $\Phi$ to ablations using only $\Phi_{\text{col}}$ or $\Phi_{\text{rot}}$.
+We find that using $\Phi$ and $\Phi_{\text{rot}}$ yields nearly identical XP performance, with $\Phi$ only slightly reducing game length.
+In contrast, $\Phi_{\text{col}}$ achieves strong SP but substantially lower XP, underperforming both $\Phi$ and $\Phi_{\text{rot}}$.
+This raises an interesting question for implementing OP in practice: How robust is OP to only partial knowledge of symmetry classes?
+We show that any single symmetry class alone may not necessarily sufficient for achieving good XP performance in the YLE.
+Alternative Neural Architectures
+Table 5: We compare several neural encoders (GATv2, GCN, and RM) and adding additional hidden layers (+1 HL, +2 HL). We find that our architecture is best among them.
+Self-Play (SP)
+Alg. V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
+SP 9C 2 $7.5\pm 0.0$.3) $89\pm 7.$ $89\pm 8.$ $92\pm 8.$ $25.0\pm 0.0$.6)
+GATv2 9C 2 $6.9\pm 0.0$.1) $87\pm 1.$ $87\pm 1.$ $100\pm 0.$ $26.0\pm 0.0$.3)
+GCN 9C 2 $4.0\pm 0.0$.2) $9\pm 8.$ $9\pm 0.$ $25\pm 21.$ $31.4\pm 0.0$.5)
+RM 9C 2 $3.0\pm 0.1$.6) $25\pm 22.$ $25\pm 22.$ $25\pm 22.$ $29.9\pm 0.1$.8)
++1HL 9C 2 $4.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0)
++2HL 9C 2 $4.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0)
+To rule out a suboptimal neural architecture, we compare our CNN–GRU policy to several graph-based encoders motivated by YLE’s relational structure: GATv2 (Brody et al., 2021), GCN (Kipf & Welling, 2016) and Relation Module (RM) (Zambaldi et al., 2018), as well as deeper MLP variants (+1HL, +2HL).
+Results in Table 5 (four seeds) show that all alternatives perform substantially worse in self-play than the CNN–GRU baseline.
+Zero-Shot Coordination Experiments with Imperfect Memory
+Table 6: We explore OP in the non-perfect memory setting Yōkai is typically played in ().
+We show that our GRU based policy is not able to learn under that setting.
+As an alternative for improved memory, we explore TrXL and show that while it improves performance, it fails to match the perfect memory setting or human players.
+For comparison, we reprint OP with perfect memory ().
+Self-Play (SP) Cross-Play (XP)
+Alg. V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
+Human 9C 2 $5.7\pm 0.0$.5) $65\pm 11.$
+OP 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
+OP 9C 2 $0.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0) $0.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0)
+OP TrXL 9C 2 $0.3\pm 0.0$.1) $1\pm 0.$ $33\pm 14.$ $1\pm 0.$ $21.7\pm 0.4$.2) $0.3\pm 0.0$.1) $1\pm 0.$ $50\pm 17.$ $1\pm 0.$ $17.8\pm 0.4$.5)
+A central aspect of Yōkai is that agents must remember the colours of moving cards.
+Our previous experiments isolated coordination by providing perfect memory; here we evaluate OP in the standard imperfect-memory setting.
+As shown in Table 6, the GRU-based OP policy fails completely without memory assistance.
+We therefore replace the GRU with Transformer-XL (TrXL) (Dai et al., 2019; Parisotto et al., 2020), an explicit memory mechanism.
+While TrXL improves performance, it remains below both the perfect-memory setting and human performance.
+This introduces an additional difficulty axis in YLE.
+Beyond increasing player count or scaling to 16C, imperfect memory substantially amplifies coordination challenges, further widening the gap between current ZSC methods and human play.
+Figure 5: Probing card colours from the hidden state in self-play and cross-play over timesteps.
+6.4 Analysing Common Ground Using Linear Probes
+To examine the belief representations learned by our agents, we train linear probes on recurrent hidden states to predict card colours at each board position using an episode-level split: for each policy, we hold out $20\%$ of self-play episodes and train on the remaining episodes.
+We evaluate on disjoint held-out SP episodes as well as XP episodes by applying each policy’s probe to that policy’s hidden states during cross-play.
+Probe accuracy serves as a linear readout of the encoded colour and thus as a diagnostic of first-order belief representations.
+As shown in Figure 5, card colours are linearly decodable above chance in SP across methods, but probe accuracy consistently decreases in XP and all remain under a perfect memory oracle.
+This reduction in decodability mirrors the observed degradation in coordination performance, suggesting that belief representations become less accessible when agents interact with unseen partners.
+Additional probing results in which we stratify by observed vs unobserved cards are shown in Appendix E and match our observations.
+7 Discussion
+YLE is a timely new ZSC benchmark.
+There is a long line of research on ZSC that almost exclusively focused on Hanabi as a benchmark (Hu et al., 2020; 2021; Cui et al., 2022; Muglich et al., 2025; Forkel et al., 2026).
+However, over the last few years, Hanabi scores have reached near optimal performance.
+Starting with OP, which produced a SP–XP gap in two-player Hanabi of $8.6\%$ ($22.07$ vs $24.06$ (Hu et al., 2020)).
+Subsequent work progressively narrowed this gap and Forkel et al.
+(2026) most recently achieved near-perfect inter-seed XP performance with a gap below $0.1\%$ ($24.47$ vs $24.48$) in two-player Hanabi, while also achieving similar performance for all other player counts.
+While this constitutes a major achievement, it also implies diminishing returns for using Hanabi as the primary stress test for ZSC.
+Saturation in Hanabi does not imply that ZSC is solved more broadly.
+While both are cooperative Dec-POMDPs, YLE differs structurally in several key respects.
+Beliefs over cards must be tracked across dynamic spatial reconfigurations, hints are ambiguous rather than guaranteed truthful, and agents must decide when sufficient common ground has been established to terminate early.
+In Hanabi, by contrast, beliefs are tied to hand slots, hints are truthful, and there is no termination decision.
+Our results in the YLE show persistent and substantial SP-XP gaps across the same families of methods.
+Even OP exhibits large degradations in both early-ending frequency and early-ending calibration in cross-play.
+In particular, XP calibration (WEE) drops close to chance level in several settings, indicating a breakdown of shared belief formation rather than merely conservative stopping behaviour.
+Importantly, the majority of our experiments focus on the simplest YLE configuration (two players, memory assistance, nine cards), indicating substantial headroom for further progress.
+Method rankings also reverse relative to Hanabi: whereas HE >> OBL >> OP in Hanabi, we observe OP >> HE >> OBL in YLE.
+This highlights the risk of overfitting to a single benchmark and motivates evaluation across diverse environments.
+YLE highlights limitations in existing ZSC methods.
+Beyond SP-XP gaps, YLE highlights structural assumptions that underlay current ZSC approaches.
+As discussed in section 5, OBL incurs substantially higher interaction costs due to longer fictitious rollouts and must operate under ambiguous hints, making belief modelling more difficult than in Hanabi.
+OP assumes that relevant symmetry classes of the Dec-POMDP are known and can be enumerated.
+In the YLE, however, symmetry structure is richer due to spatial layout and card movement, and multiple interacting symmetry sources exist.
+Enumerating all such symmetries is difficult in practice, and, as we show, enforcing invariance to only a subset may leave residual symmetry-breaking conventions.
+Finally, while HE effectively coordinates in Hanabi, increasing entropy in YLE often destabilises the learned policy before optimal cross-play performance is reached, particularly in multi-player settings.
+Taken together, these results indicate that techniques successful in Hanabi do not directly transfer to YLE.
+The reversal in method rankings further suggests that apparent ZSC progress may be benchmark-specific, underscoring the need for structurally diverse evaluation environments.
 8 Related Work
-Current benchmarks study ToM reasoning in AI mostly from a static observer’s perspective Baker et al.
-(2011, 2017); Rabinowitz et al.
-(2018); Bara et al.
-(2021, 2023); Fan et al.
-(2021); Duan et al.
-(2022); Kim et al.
-(2023); Bortoletto et al.
-(2024c); Gandhi et al.
-(2023), framed as a supervised learning task – an approach with notable limitations Aru et al.
-(2023); Bortoletto et al.
-(2024a).
-Other benchmarks evaluated agents in interactive environments, such as the Hanabi Learning Environment (HLE) Bard et al.
-(2020) and SymmToM Sclar et al.
-(2022).
-In contrast to Hanabi, Yōkai requires spatial-temporal belief reasoning, requiring agents to bind beliefs to moving cards Fernandez et al.
-(2023).
-Compared to SymmToM, YLE offers greater strategic depth and stronger pressure to develop ToM.
-In SymmToM, heuristics outperform all baselines and communication is unrestricted, while in YLE, misused hints can irreversibly lock cards in losing positions.
-Unlike Hanabi and SymmToM, YLE also introduces a high-stakes belief-based decision: agents may end the game early for a higher payoff, sacrificing future information and relying more on their beliefs.
-A second line of work investigated whether neural networks represent mental states internally Bortoletto et al.
-(2024a); Zhu et al.
-(2024).
+Current benchmarks study ToM reasoning in AI primarily from a static observer’s perspective (Baker et al., 2011; 2017; Rabinowitz et al., 2018; Bara et al., 2021; 2023; Fan et al., 2021; Duan et al., 2022; Kim et al., 2023; Bortoletto et al., 2024b; Gandhi et al., 2023; Bortoletto et al., 2025), often framed as supervised learning problems (Aru et al., 2023; Bortoletto et al., 2024a).
+In contrast, interactive benchmarks such as the Hanabi (Bard et al., 2020) and SymmToM (Sclar et al., 2022) evaluate ToM in multi-agent interactive settings.
+YLE extends interactive ToM benchmarks by introducing dynamic entity movement, ambiguous communication, and early termination decisions within a cooperative Dec-POMDP.
+Yōkai was first explored in AI research in the context of formal logics (Fernandez et al., 2023), and, additionally, has been explored in a preliminary study on the feasibility of training self-play RL agents for Yōkai (Yang, 2024).
+The present work instead introduces the YLE as an open-source benchmark for belief tracking and zero-shot coordination.
+A second line of work investigates whether neural networks internally represent mental states (Bortoletto et al., 2024a; Zhu et al., 2024).
 For example, Matiisen et al.
-Matiisen et al.
-(2022) probed DRL agents for intention representations in cooperative navigation.
-We extend probing to YLE, where beliefs evolve dynamically and asymmetric knowledge emerges over time – showing the value of probing in a more complex environment.
-Finally, YLE contributes to the broader family of cooperative multi-agent RL (MARL) benchmarks Samvelyan et al.
-(2019); Carroll et al.
-(2019); Bard et al.
-(2020); Ruhdorfer et al.
-(2024), but with a strong emphasis on ToM and common ground reasoning.
-While many techniques were developed for Hanabi, our results indicate they do not transfer directly to the YLE, thus highlighting the new challenges posed by our benchmark and requiring new modelling approaches.
-7 Conclusion & Future Work
-We introduced the Yōkai Learning Environment – a multi-agent RL benchmark where effective collaboration requires belief tracking over space and time.
+(2022) probe RL agents for intention representations in cooperative navigation.
+We extend probing analyses to YLE, where hidden representations must encode evolving beliefs under partial observability.
+Alongside YLE, OvercookedV2 (Gessler et al., 2025) has recently been proposed as a benchmark for ZSC, augmenting the Overcooked environment (Carroll et al., 2019) with partial observability and asymmetric information.
+YLE differs in that it introduces substantially larger state and action spaces and, in configurations with more than three players or 16 Yōkai cards, poses a significant challenge even in SP.
+Additionally, YLE offers a full reference implementation of OBL, which previously achieved state-of-the-art inter-seed cross-play performance in the HLE.
+Our results show that methods with strong performance in Hanabi perform substantially worse in YLE and that algorithm rankings change, highlighting the importance of diverse benchmarks for evaluating ZSC algorithms.
+YLE also contributes to the broader family of cooperative multi-agent RL benchmarks (Samvelyan et al., 2019; Ruhdorfer et al., 2025b; Tomilin et al., 2025), with a particular emphasis on belief tracking, ToM, and common ground formation.
+Finally, we consider the ZSC setting as introduced by Hu et al.
+(2020) and further clarified in (Treutlein et al., 2021).
+Other broader notions of coordination with novel partners exist and are often studied under the label of ad-hoc teamwork (AHT) (Stone et al., 2010).
+Representative AHT methods are typically population- and diversity-based and include (Strouse et al., 2021; Lucas & Allen, 2022; Xue et al., 2022; Zhao et al., 2023; Yan et al., 2023; Hui et al., 2025; Ruhdorfer et al., 2025a).
+In this work, we focus on ZSC first to isolate the coordination challenges posed by the YLE, reasoning that AHT progress on the YLE is only possible after making progress on the ZSC problem.
+We thus view evaluating AHT methods in the YLE as valuable future work after ZSC performance has improved, similar to the progression observed in Hanabi research (Dizdarevic et al., 2025).
+9 Conclusion
+We introduced the Yōkai Learning Environment – a new multi-agent RL benchmark in which effective collaboration requires tracking and maintaining beliefs over space and time.
 Implemented in JAX, YLE supports end-to-end GPU training at hundreds of thousands of steps per second.
-We show that RL agents on the YLE struggle to retain knowledge, to generalise across partners, or to identify when coordination has succeeded – a high-stakes decision reflected in their inability to end games early.
-YLE thus offers a scalable, diagnostic testbed for advancing common ground reasoning in collaborative AI.
-While our work focused on evaluating RL agents, future work might use this benchmark to evaluate other systems, like language agents Xi et al.
-(2023); Wang et al.
-(2024).
-We envision that Yōkai can play a major role in isolating and assessing critical aspects of these systems that realte to memory, spatial reasoning and common ground tracking.
-Acknowledgements
+Empirically, we show that state-of-the-art ZSC methods that achieve near-perfect inter-seed cross-play in Hanabi exhibit persistent SP-XP gaps in YLE, including degraded early-ending frequency and calibration.
+These results indicate that coordination in YLE places additional structural demands on belief formation and common ground reasoning and that apparent ZSC progress may be benchmark-specific.
+The YLE thus provides a complementary benchmark for evaluating ZSC beyond current environments.
+Broader Impact Statement
+This work introduces a new environment for studying fundamental aspects of cooperation in reinforcement learning agents.
+As such, it supports basic research and is currently far removed from direct societal deployment.
+However, because our goal is to improve human–AI collaboration, caution is warranted when developing systems.
+Any future applications involving real users must prioritise safety, transparency, and accountability.
+Acknowledgments
 The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting C.
@@
 Shi for numerous insightful discussions.
-Appendix A Appendix: Infrastructure & Tools
-Servers and software
-We ran our experiments on two different server systems.
-Both are running Ubuntu 24.04.
-One is equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs.
-The other is equipped with NVIDIA H100-NVL GPUs with 94GB of memory and AMD EPYC 9454 CPUs.
+J.
+Foerster is partially funded and J.
+Forkel is funded by the UKRI grant EP/Y028481/1 (originally selected for funding by the ERC).
+Supplementary Materials
+The following content was not necessarily subject to peer review.
+Appendix A Servers and Software
+We ran our experiments on server systems running Ubuntu 24.04 equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory, NVIDIA H100-NVL GPUs with 94GB of memory and a third server system featuring NVIDIA L40S with 48 GB of memory.
 All training runs are executed on a single GPU only.
-We trained our models using JAX Bradbury et al.
-[2018], Flax Heek et al.
-[2023] and Optax DeepMind et al.
-[2020].
-In terms of tools, we did our analysis using NumPy Harris et al.
-[2020], Pandas Team [2020], SciPy Virtanen et al.
-[2020] and Matplotlib Hunter [2007].
-Our TransformerXL implementation is based on the one provided by Hamon [2024].
-Our single-file implementation of MAPPO is based on the one provided by JaxMARL Rutherford et al.
-[2023] with additional inspiration taken from Hamon [2024].
-Accesability and reusability
+We trained our models using JAX (Bradbury et al., 2018), Flax (Heek et al., 2023) and Optax (DeepMind et al., 2020).
+In terms of tools, we did our analysis using NumPy (Harris et al., 2020), Pandas (Team, 2020), SciPy (Virtanen et al., 2020) and Matplotlib (Hunter, 2007).
+Our TransformerXL implementation is based on the one provided by Hamon (2024).
+Our single-file implementation of MAPPO is based on the one provided by JaxMARL (Rutherford et al., 2023) with additional inspiration taken from Hamon (2024).
+Appendix B Accessibility and reusability
 The Yōkai Learning Environment (YLE) integrates directly into JaxMARL and implements its standard environment interface.
 As a result, any baseline method implementation of JaxMARL can be adapted to our environment quickly and with minimal effort.
+Upon acceptance, we will release the full environment implementation, training code for all reported baselines (including OP, HE-IPPO, and OBL), configuration files for reproducing the reported experiments, probing scripts, and pretrained model checkpoints.
+We have coordinated with the JaxMARL authors Rutherford et al.
+(2023), and upon acceptance YLE will be integrated into the official JaxMARL repository.
 This makes our open-source environment easily accessible to future researchers who will benefit from our reference implementations and those provided by the wider community.
@@
 The project is accessible under https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env.
-Training times and compute requirements
-Training times vary between server setup and agent.
-Our longest training agent is ToMTrack-OP+C+R since it needs to compute memory updates for the TransformerXL and the asymmetric domain randomisation used in Other-Play.
-Experiments with ToMTrack-OP+C+R take 8 hours to complete per seed on our H100 setup and around 24 hours on our V100 setup.
-The CNN-M baselines train quickest:  6 hours on the H100 server and  14 hours on a V100 server.
-All our experiments are runnable on a single V100 with 32GB VRAM.
-Appendix B Fast, Scalable, and Parallel Yōkai in Graph-based JAX
+Appendix C Fast, Scalable, and Parallel Yōkai in Graph-based JAX
 We first describe in detail how our environment operates and then share additional details on reinforcement learning in the YLE.
-B.1 JAX-based Implementation
+C.1 JAX-based Implementation
 We carefully engineer our environment to be end-to-end usable on a GPU directly and make it just-in-time compilable using JAX.
@@
 […]
-Resulting Performance
-Num Envs $n$ 512 1,024 2,048
-2-player 3x3 139,107 274,669 559,251
-2-player 4x4 135,662 144,636 148,269
-Table 3: We evaluate steps-per-second (SPS) of our environment on a single Nvidia H100 by taking random actions in $n$ parallel environments. Our environment scales to hundreds of thousands of SPS.
-We present how our environment performs in terms of steps-per-second across many parallel environments.
-All experiments were run by taking 1000 random consecutive actions in $n$ parallel environments.
-Table 3 shows results.
-Our environment scales similarly to other recently proposed JAX-based RL environments, e.g.
-Rutherford et al.
-[2023]; Nikulin et al.
-[2023]; Ruhdorfer et al.
-[2024]; Matthews et al.
-[2024].
-B.2 Reinforcement Learning in the YLE
+C.2 Reinforcement Learning in the YLE
 We want to continue discussing additional aspects of reinforcement learning in the YLE.
@@
 Figure 6: Examples of start, winning, and illegal card configurations for a Yōkai game with nine cards and three colours.
-While in the original game, cards can be placed in any legal configuration, our environment limits the configuration to be in a predefined field of size $g\times g$ where $g=9$ in the 3x3 and $g=10$ in the 4x4 version.
+While in the original game, cards can be placed in any legal configuration, our environment limits the configuration to be in a predefined field of size $g\times g$ where $g=9$ in the 9C and $g=10$ in the 16C version.
 This enables the efficient computation of legal moves but prohibits some rare card configurations.
 While computing the exact number of disallowed configurations is infeasible, we find them to be extremely uncommon in practice.
-For example, let’s assume a 2-player 3x3 game on a 9x9 grid.
+For example, let’s assume a 2-player 9C game on a 9x9 grid.
 In such a setting, there a 4 hints and thus a total of 8 move actions for both players together.
@@
 We find that policies rarely to never use to border region of the board.
-Our best-performing policy ToMTrack uses the border $0.0\pm 0.0\%$ of the time when evaluated over all three seeds in $1000$ games each.
-We also discuss this in section 5.7.
 The action space in detail
@@
 $N_{a}=1+|Y|+N_{moves}+|H|+N_{hint\_moves}+1.$ (9)
-For example, the action space is of size $1,068$ for 2-player 3x3 games and $2,322$ for the 4x4 version.
+For example, the action space is of size $1,068$ for 2-player 9C games and $2,322$ for the 16C version.
 YLE is a symmetric ToM environment
 Sclar et al.
-[2022] defines symmetric ToM environments as environments that (1) have a symmetric action space, (2) feature imperfect information, (3) allow the observation of others, and (4) require information-seeking behaviour.
+(2022) defines symmetric ToM environments as environments that (1) have a symmetric action space, (2) feature imperfect information, (3) allow the observation of others, and (4) require information-seeking behaviour.
 YLE features all four and thus qualifies.
@@
 Number of Players, Hints and the Difficulty
-4x4 3x3
+Table 7: Number of hint cards according to the number of players and the environment version. The 9C version features three colours and thus has no three-colour hints.
+16C 9C
 Number of Players 2 3 4 2 3 4
@@
 Sum 7 9 10 4 5 6
-Table 4: Number of hint cards according to the number of players and the environment version. The 3x3 version features three colours and thus has no three-colour hints.
 How many and which hint cards are available depend on the environment version and the number of players.
@@
 Instead, the number of cards one player can directly observe themself decreases with the number of players.
-For 4x4 YLE in the 2-player version, any agent can observe up to 14 of the 16 cars themselves (8 of 9 in 3x3 Yōkai).
-This shrinks to 10 of 16 for the 4x4 version and 6 of 9 for the 3x3 version when 4 players take part.
+For 16C YLE in the 2-player version, any agent can observe up to 14 of the 16 cars themselves (8 of 9 in 9C Yōkai).
+This shrinks to 10 of 16 for the 16C version and 6 of 9 for the 9C version when 4 players take part.
 Note how in the 3- and 4-player versions, agents are thus required to both work with less first-hand information and need to rely more on hints and interpreting the moves made by those around them.
@@
 This will usually highly restrict the agents’ ability to sort cards.
-B.3 Second Order ToM Reasoning Example
-Figure 12: A second-order ToM reasoning example over multiple timesteps in 2-player YLE.
-Refer to caption beliefs that Refer to caption beliefs that he (Refer to caption ) knew that cards 1 and 2 are of the same colour.
+C.3 Second Order ToM Reasoning Example Extended
+Figure 8: Here we reexplain the second-order ToM reasoning example from the main paper over multiple timesteps in 2-player YLE in additional detail (see text): Refer to caption beliefs that Refer to caption beliefs that he (Refer to caption ) knew that cards 1 and 2 are of the same colour.
 Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground.
 In the future, they can now both observe and move different cards to finish early as they have one less card that needs to be observed.
-Figure 12 displays an example for second-order ToM reasoning in 2-player Yokai.
-It shows an agent ([Uncaptioned image] ) making inferences about what they assume the other agents ([Uncaptioned image] ) knowledge is due to their own previous action.
+Figure 8 displays the same example for second-order ToM reasoning in 2-player Yokai as the main paper.
+Here, we expand on the example.
+Figure 8 shows an agent ([Uncaptioned image] ) making inferences about what they assume the other agents ([Uncaptioned image] ) knowledge is due to their own previous action.
 In the example [Uncaptioned image] infers the colour of the card due to the behaviour of the other agent.
@@
 Note that such higher-order reasoning steps are even likelier to occur in settings involving more partner agents.
-Appendix C Training Details
+Appendix D Training Details
 We share all the details of the training here, including all hyperparameters used in training and all details on the network architectures and how we tuned them.
 In general, we have tuned many aspects of our policy to arrive at the strongest agents we could for our experiments, see below.
-Finally, the number of parameters per model is shown in Table 6.
 We took extra care in making these equal to make our results comparable.
-C.1 Reinforcement Learning Details
+D.1 Reinforcement Learning Details
+Table 8: Hyperparameters of the learning process.
 Description Value
-Optimiser Adam Kingma and Ba [2015]
+Optimiser Adam (Kingma & Ba, 2015)
 Adam $\beta_{1}$ $0.9$
@@
 Reward Shaping Yes (linearly annealed)
-Table 5: Hyperparameters of the learning process.
-Our work employs a standard MAPPO algorithm Yu et al.
-[2022] for all experiments.
-MAPPO is an extension of PPO Schulman et al.
-[2017] to the multi-agent setting.
-It separates the actor and the critic network.
-Since the actor and critic are separated and the critic is discarded after training, MAPPO allows giving the critic access to privileged information during training.
-This leads to the good performance of MAPPO in cooperative multi-agent tasks, compare Yu et al.
-[2022], which is why we choose it in our work.
+Our work employs a standard IPPO algorithm (de Witt et al., 2020) for all experiments.
 We present all parameters of the learning process in Table 8.
-C.2 Neural Network Architectures
-Model # Parameters
-GCN-M 317,048
-GATv2-M 317,560
-RM-M 354,168
-CNN-M 322,156
-CNN-LSTM 470,252
-CNN-GRU 437,484
-CNN-S5 388,524
-CNN-TrXL 400,940
-ToMTrack 404423
-ToMTrack-SAD 417863
-Table 6: Number of parameters of all models used in this work.
-Our work employs several encoder architectures.
-We present all the details on them below.
-In general, we aimed to make comparisons fair by giving all compared models a similar number of parameters, also see Table 6.
-The biggest advantage in terms of parameter size was RM-M for the spatial reasoning and CNN-LSTM for the temporal reasoning experiment, but neither performed the best.
-C.2.1 General Policy Architecture
+Only for OBL we reduce the number of parallel environments to $256$ since with $K=8$ we are otherwise unable to fit the OBL policies into memory.
+Note that this actually increases the number of gradient updates OBL policies recieve.
+D.2 Neural Network Architectures
+Our work employs several encoder architectures. We present all the details on them below.
+General Policy Architecture
 In general, in our work, no matter the encoder, the final policy network always has the same structure.
-That is given an embedding $e_{\text{encoder}}$ produced by an encoder network $f_{\text{encoder}}$ and possibly a memory architecture, we pass it to a stack of dense layers and either a value or action head.
-We always use $4$ dense layers before the heads with ReLU Fukushima [1975]; Nair and Hinton [2010] as the activation function between the layers and a hidden dimension of $128$.
+That is given an embedding $e_{\text{encoder}}$ produced by an encoder network $f_{\text{encoder}}$, we pass the embedding to a linear projection, then a GRU (Cho et al., 2014) and finally either a to value or action head.
+We use ReLU (Fukushima, 1975; Nair & Hinton, 2010) as the activation function.
 The layers are initialised using orthogonal initialisation with a scale of $\sqrt{2}$ and bias 0.
 We have determined the number of policy layers and their size using hyperparameter tuning.
-We compared $2$, 33 and $4$ layers with hidden dimension $32$, $64$ and $128$.
-We found that the largest configuration outperforms the others.
 Additionally, for all neural network architectures detailed below, we also varied and explored: The number of encoder blocks (2, 3 and 4), the hidden dimension of the encoder blocks, the aggregation function after the encoder (mean, sum, multiply and concatenate) and whether to anneal the dense rewards (False and True).
+We also compared addition additional hidden layers between GRU and the action/value heads (0, +1HL, +2HL).
 Finally, the value head produces a singular scalar output and is initialised using orthogonal initialisation with a scale of $0.00$ instead of $\sqrt{2}$.
 We initialise the action head in the same way but it produces logits for all possible actions instead.
-The critic network is always parameterised the same as the actor network, but the input is different.
-Since we are using MAPPO our critic receives as input a so-called world state that combines the individual observations of all agents.
-This is possible in MAPPO due to centralised training and decentralised execution Foerster et al.
-[2016].
-As the world state, we concatenate the individual observations along the feature dimension.
-C.2.2 GCN, GATv2 and the pre-embedding of Nodes
+GCN, GATv2 and the pre-embedding of Nodes
 Figure 9: A sketch of the GNN and pre-embedding architecture used in this work.
-We use two graph neural networks in our work, one based on graph convolutions Kipf and Welling [2016] and one based on the attention mechanism Brody et al.
-[2021].
+We use two graph neural networks in our work, one based on graph convolutions Kipf & Welling (2016) and one based on the attention mechanism Brody et al.
+(2021).
 We pick these two as they stem from two different families of graph learning architectures and they thus are representative of a larger body of work.
@@
 First, both graph architectures observe an input graph that features a virtual node, similar to the hub-nodes of Abdessaied et al.
-[2024] or the [VNode] token of Ying et al.
-[2021].
+(2024) or the [VNode] token of Ying et al.
+(2021).
 A virtual node is an empty node that connects to all other nodes in the graph.
@@
 Second, our work uses pre-embeddings of node-feature modalities as described in Bortoletto et al.
-[2024d].
+(2024c).
 We take the node features of size $n\times f$ and embed them per modality via a separate dense encoder before then concatenating all of the resulting vectors into new node features shaped $n\times f^{\prime}$.
@@
 see Bortoletto et al.
-[2024d].
+(2024c).
 Specifically, for each modality $m\in M$ we train a dense layer $f_{m}$ that embeds the modality per node into an embedding $e_{m}\in\mathbb{R}^{b\times t\times n\times 2}$ for the current player embedding or $e_{m}\in\mathbb{R}^{b\times t\times n\times 4}$ for all other modalities where $b$ is the batch size, $t$ the time and $n$ the number of nodes.
@@
 The resulting embedding is then concatenated along the node dimension and fed to the rest of the policy.
-C.2.3 Relation Module
+Relation Module
 Our work adapts the relation module of Zambaldi et al.
-[2018].
+(2018).
 We only use the self-attention mechanism of their proposed relation module and skip the CNN encoder they propose alongside the self-attention layers.
 This is because our setting deals with an environment where agents observe the position of cards directly as opposed to the setting of Zambaldi et al.
-[2018] that deals with image inputs directly.
+(2018) that deals with image inputs directly.
 Thus, in our relation module, non-local interactions between all cards are simply computed using self-attention directly.
@@
 Our relational module is also applied $4$ times, features $4$ heads, and its output dimension per layer is $64$.
-C.2.4 CNN
+CNN
 Our work compares a CNN encoder and determines it to be the best-performing model.
 As described earlier, the YLE returns an image-like observation to be processed by our CNN.
-The exact size of the input tensor to the CNN depends on the YLE settings, but in 2-player 3x3 YLE, the input dimension is $9\times 10\times 13$, which we will use as a running example.
+The exact size of the input tensor to the CNN depends on the YLE settings, but in 2-player 9C YLE, the input dimension is $9\times 10\times 13$, which we will use as a running example.
 It includes the same features as discussed above, with some minor changes.
@@
 Our convolutional stack uses $4$ 2d convolutions with $64$ filters each and ReLU applied between them.
-The kernel size is $(3,3)$, we apply a stride $1$ and valid padding.
+The kernel size is $(3,3)$, we apply a stride 11 and valid padding.
 The resulting embedding is flattened to produce the final embedding to be passed on to later layers.
-C.2.5 LSTM and GRU
-Next to the encoder architectures, our work compares several memory architectures.
-Here, memory architecture is used as a term to denote classical RNNs as well as structured state space or transformer Vaswani et al.
-[2017] architectures.
-LSTMs Hochreiter and Schmidhuber [1997] and GRUs Cho et al.
-[2014] are commonly used in RL and such are natural first approaches for our environment.
-They provide an implicit memory through their hidden state that is computed at every step from the current observation encoding and the previous hidden state.
-In both cases, we take the embedding generated by the encoder, $e_{\text{encoder}}$, pass it through a dense layer $f_{\text{pre\_rnn}}$ before finally applying either the LSTM or GRU on the resulting embedding.
-$f_{\text{pre\_rnn}}$ embeds $e_{\text{encoder}}$ to the hidden dimension of the RNN.
-As the hidden dimension of the RNN we pick $128$ in our work.
-We also experimented with using $256$ but found no difference in performance.
-We use orthogonal initialisation for both the GRU and the LSTM.
-C.2.6 S5
-As already described, we use S5 Smith et al.
-[2023] layers because of the great performance of structured state-space models on memory tasks in deep model-based RL Samsami et al.
-[2024].
-Similar to LSTM and GRU, S5 only provides an implicit memory mechanism.
-We use a similar setup as with both the LSTM and GRU layers, i.e.
-we also use a $f_{\text{pre\_s5}}$ embedding layer that embeds $e_{\text{encoder}}$ to the hidden dimension of our S5 layer.
-We use $2$ S5 blocks, $1$ layer, place the LayerNorm Ba et al.
-[2016] before the S5 layer and use the activation function proposed in the original work Smith et al.
-[2023].
-As before we use $128$ as the hidden dimension of S5.
-C.2.7 TransformerXL
-Finally, as already introduced TransformerXL Dai et al.
-[2019] is an explicit memory mechanism that shows great performance in memory-intensive RL tasks when used correctly Parisotto et al.
-[2020].
+TransformerXL
+Next to the encoder architectures, our work compares memory architectures.
+TransformerXL (Dai et al., 2019) is an explicit memory mechanism that shows great performance in memory-intensive RL tasks (Parisotto et al., 2020).
 Our usage strictly follows the one described by Parisotto et al.
-[2020] as implemented by Hamon [2024].
-Similar to the other memory architectures, we also first pass the embedding received by the encoder through a dense layer $f_{\text{pre\_trxl}}$ before handing the resulting embedding to the transformer.
-While LSTM, GRU and S5 use $128$ as their hidden dimension, we only use a hidden dimension of $64$ for both the dense layers within the transformer as well as the projection layers in multi-head attention.
+(2020) as implemented by Hamon (2024).
+Similar to our GRU architecture, we first pass the embedding received by the encoder through a dense layer $f_{\text{pre\_trxl}}$ before handing the resulting embedding to the transformer.
+We use a hidden dimension of $64$ for both the dense layers within the transformer as well as the projection layers in multi-head attention.
 This is to keep the number of parameters consistent between the architectures.
@@
 Finally, we set the length memory window to $32$.
-$32$ is long enough to recall all of the previous moves in the 3x3 YLE.
-Memories from a past game – as marked by the done signal – are masked.
+$32$ is long enough to recall all of the previous moves in the 9C YLE.
+Memories from a past game are masked.
 We apply standard positional encoding based on sinus waves.
-C.3 ToMTrack
-Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
-ToMTrack w/o $f_{\text{belief}}$ explicit $0.9\pm 0.0$.5) $31.2\pm 1.4$.8) $1.8\pm 0.0$.4) $0.0\pm 0.0$.0)
-ToMTrack w/o action explicit $1.4\pm 0.0$.4) $41.2\pm 1.1$.7) $2.0\pm 0.0$.3) $0.0\pm 0.0$.0)
-ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
-Table 7: ToMTrack ablations. We evaluate ToMTrack without the additional belief representation learning and the additional action encoding and find that only both together improve performance.
-ToMTrack is a direct extension of the CNN-TrXL baselines and shares all the parameters mentioned earlier with both the CNN as well as the TransformerXL work.
-There are two extensions we add to arrive at ToMTrack.
-The first one – the auxiliary belief prediction task – is already mostly described in the main text.
-To generate the distribution over possible card colours, we add a colour prediction head $f_{\text{belief}}$ to our policy.
-$f_{\text{belief}}$ takes as input the embedding generated by the stack of dense layers and produces an output $e_{\text{belief}}\in\mathbb{R}^{b\times t\times|Y|c}$ that resembles the prediction over all cards and colours.
-For the 3x3 version $|Y|c=9\cdot 3=27$.
-To obtain the final predictions, we reshape $e_{\text{belief}}$ to be of shape $b\times t\times|Y|\times c$ and apply the Softmax function of the last dimension to get a prediction per card.
-The method only has one hyperparameter $\lambda$ – which balances the losses - and we pick $\lambda=0.1$ in this work.
-The second extension – the active action decoding – is realised via an additional input stream.
-Before passing the previous action to this stream, we first decode it into a reduced dimension that serves as a dense representation of that action of shape $4+|Y|+2+|H|+1$.
-The first $4$ bits encode the move type alias the substep.
-The next $|Y|$ bits encode the Yōkai card affected by the current action, if no card is affected, all are set to 0.
-The next $2$ bits encode the position affected by the action.
-For instance, if the previous active action was a move action, the resulting coordinates are encoded there.
-The next $|H|$ bits encode whether and if so which hint card is used in an action, and finally, the last bit encodes whether the previous action was to terminate the game.
-For a 2-player 3x3 YLE game $4+|Y|+2+|H|+1=20$.
-Note that this is a much denser representation compared to just encoding the raw number.
-The resulting encoding $e_{\text{action}}$ is processed by its own dense layer and then concatenated with the output of the encoder before being passed to the rest of the policy network.
-Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
-C.4 Other-Play in Yōkai
+D.3 Additional Details for Other-Play in Yōkai
 Figure 10: An example of how both agents act in rotated and recoloured environments to break symmetries for zero-shot coordination.
@@
 OP presents agents with a special asymmetric domain randomisation that makes it impossible for agents to coordinate on breaking symmetries in certain ways.
-In Hanabi, this is achieved by a recolouring operation that shows each agent a different recolouring of the game state Hu et al.
-[2020].
+In Hanabi, this is achieved by a recolouring operation that shows each agent a different recolouring of the game state (Hu et al., 2020).
 Therefore, to train agents via OP for the YLE, one first needs to find all classes of symmetries that exist in an environment, and then, second, find an appropriate domain randomisation operation.
@@
 Note that these are not the only ways for agents to learn to break symmetries.
-Recent work Treutlein et al.
-[2021] showed that agents also might coordinate to break symmetries in arbitrary ways over multiple timesteps.
+Recent work (Treutlein et al., 2021) showed that agents also might coordinate to break symmetries in arbitrary ways over multiple timesteps.
 It is impossible to list all possible classes of symmetries for an environment with enough complexity.
@@
 In our work, we thus focus on showing that while Hanabi only requires breaking one symmetry for efficient ZSC while the YLE – due to its additional complexities – requires additional effort to improve ZSC performance, which is still far from optimal.
-Appendix D Additional Probing Details
-Over the duration of $1000$ games, we collect a probing dataset $\mathcal{D}=\{x^{(i)},y^{(i)}\}$.
+Appendix E Additional Probing Details
+Figure 11: Probing results where only known card colours from openend cards where probed.
+We can see that agents start of representing card colours in a linearly decodeable way but as the game progresses and more cards are opened, accuracy decreases.
+Figure 12: Probing results where only unknown card colours from unopenend cards where probed.
+As the game progresses card colours become increasingly linearly decodable but remain under a perfect baseline.
+Only OBL and OP represent unknown cards in a linear-decodable fashion above random chance.
+Over the duration of $5,000$ games, we collect a probing dataset $\mathcal{D}=\{x^{(i)},y^{(i)}\}$.
 The probing datasets map the activations of an agent to a label $z$.
 A probe then is a function $g:f_{l}(x)\to\hat{z}$ and trained on $\mathcal{D}$ where $f$ is a neural network and $f_{l}(x)$ is the intermediate representation of $x$ at layer $l$.
-In our case, we train simple logistic regressions.
 We always probe before the action head of our agent.
 As highlighted in the main text, during all experiments, we train one and evaluate it on a hold-out test set $\mathcal{D}_{\text{test}}\subset\mathcal{D}$.
-Additionally, note that many of these tasks feature imbalanced datasets since the does-not-know label is far more likely to occur than any other label.
-We correct for that by weighing the loss per class and subsampling the does-not-know label to the largest other class.
-During all experiments, we train one probe per card – 9 in total – and average the scores computed on a hold-out test over all 9.
-Appendix E Additional Experimental Results
-E.1 Additional Training Curves
-Figure 15: Return, success rate and numbers of clusters over time per encoder while given access to perfect memory.
-Figure 16: Return, success rate and numbers of clusters over time per memory architecture.
-Figure 17: Return, success rate and numbers of clusters over time for ToMTrack, ToMTrack-SAD, CNN-TrXL and the CNN-M architecture.
-For completeness, we show how the most important metrics of our agents developed over time.
-Figure 15 shows the curves of the spatial-reasoning experiment, Figure 16 of the memory-reasoning experiment and lastly Figure 17 reports the same for ToMTrack.
-E.2 Additional Zero-Shot Coordination Metrics
-Figure 18: Zero-shot coordination performance measured by success-rate, average return and number of cards clustered.
-SP and OP+C are consistently outperformed by OP+C+R, showing that beyond cards’ colour, agents need to focus on other reasoning types.
-Figure 19: We show training curve metrics for our agents trained with other-play to demonstrate convergence.
-Recall that these metrics measure self-play performance rather than ZSC.
-ZSC metrics are shown in Figure 6 and Figure 18.
-In Figure 18 we show additional metrics for the ZSC evaluation in our main body of work. Training curves for our other-play agents are shown in Figure 19 to show convergence behaviour.
-E.3 2- versus 4-player Games
-Figure 20: ToMTrack in the two- and four-player 3x3 YLE.
-In four-player YLE more hints are available to be placed, and as such, we observe higher total rewards collected by the ToMTrack-4P agent, even though it wins fewer games and clusters fewer colours.
-In Figure 20 we show how two- and four-player settings compare for our best model, ToMTrack.
+During all experiments, we train one probe per position (81 total, 12 seeds = 972 probes) and average the scores computed on a hold-out test over all.
+In the main text, we experiment with probing all cards (no matter if they have been observed).
+In Figure 11 and Figure 12 we show the opened vs.
+unopenend cards distinction.
 Appendix F Human Performance in Yōkai
@@
 To make results as comparable as possible, humans were instructed to avoid any form of non-verbal communication.
+Each pair played five evaluation rounds, and we report mean return across games.
+Early termination was allowed according to the same rules as in the agent environment.
 The study was conducted under an umbrella agreement of the relevant institutional review board.
@@
 Appendix G Example Games
-Figure 21: We show an example game in which ToMTrack wins in self-play.
+Figure 13: We show an example game in which an agent wins in self-play.
 For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom.
@@
 In turns 1 and 3, for instance, it gains knowledge of two separate green cards but only groups them in turn 5 after agent 1 places a corresponding gbgb-hint in turn 4.
-Figure 22: We show an example game in which ToMTrack loses in self-play.
+Figure 14: We show an example game in which an agent loses in self-play.
 For each turn, we display the ground truth information at the top and the current observable information – essentially the agents’ observation – at the bottom.
@@
 Additionally, note that in turns 4 and 6, agent 1 establishes the identity of all red cards but fails to recall that in turn 6, therefore also misplaying.
-We show an example game in which ToMTrack wins in self-play in Figure 21 and one where it loses in Figure 22.
-Appendix H Details on Other Attempts
-While ToMTrack provides a mechanism for learning representations that represent card colours, it does not necessarily encourage the agents to maximise their shared or individual knowledge.
-Consequently, we were inspired to also experiment with a variant of ToMTrack that is not trained via an auxiliary task but rather uses $\mathcal{L_{\text{CE}}}$ as intrinsic motivation for both agents; ToMTrack-I.
-ToMTrack-I agents are encouraged to form an implicit belief model that aligns well with the true state of the environment by adding an intrinsic reward $r_{i}$ to the reward based on the negative cross-entropy loss of the actor classification head:
-$r=r+\lambda_{\text{s}}*r_{\text{s}}+r_{\text{i}}\text{, where }r_{\text{i}}=-\mathcal{L}^{\text{CE}}_{a}.$ (9)
-In this formulation, the agents thereby receive a higher reward when the belief model’s prediction loss $\mathcal{L}^{\text{CE}}_{a}$ is low.
-This approach is loosely inspired by Oguntola et al.
-[2023].
-We originally assumed that this would achieve superior results but found that the intrinsic reward confused the agents to a point where they barely learned a useful policy altogether, falling below baseline performance (i.e.
-CNN-TrXL).
-Future work should study more sophisticated ideas based on this concept.
-Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
-ToMTrack (SAD) explicit $0.7\pm 0.0$.4) $21.5\pm 1.3$.4) $1.5\pm 0.0$.4) $0.0\pm 0.0$.0)
-Table 8: Additional ToMTrack-SAD results. This approach underperforms our baselines from Table 1 considerably and was thus excluded from the main paper.
-Previous work on Hanabi has proposed the simplified action decoder (SAD) Hu and Foerster [2020].
-SAD agents can take two actions during training that other agents can observe: a normal environment action and a greedy action.
-Other agents are allowed to observe both.
-The intuition is that during exploration, the true intentions of agents are blurred with explorative actions and thus less informative for others.
-Hu and Foerster [2020] implements their idea on top of epsilon-greedy Q-learning.
-Since during evaluation $\epsilon=0$, all actions are greedy and are given as input to both action inputs heads of other agents.
-During evaluation, $\epsilon=0$ and both actions are greedy.
-SAD showed improved performance in Hanabi, including the Bayesian action decoder Foerster et al.
-[2019].
-We have adapted this idea as ToMTrack-SAD to our policy learning approach.
-Since we do not have an $\epsilon$ to set, we adopt their idea by also allowing agents to take two actions during training: One sampled from the agents’ produced action distribution is used as the environment action that gets executed, and one greedy action is determined as the legal action with the highest associated logit.
-The true environment action is duplicated during evaluation with similar reasoning.
-ToMTrack-SAD is our extension of ToMTrack.
-In ToMTrack-SAD both actions are encoded as described earlier and concatenated before being fed to the same dense layer.
-Since we use MAPPO and no Q-learning approach as in the original work Hu and Foerster [2020], our agents pick their normal environment action given their current set of parameters $\theta_{a}$ in state $s_{t}$ via
-$a_{\text{env}}\sim\pi(\cdot|s_{t},\theta_{a})$ (10)
-and their second greedy action as
-$a_{\text{greedy}}=\arg\max_{i}\pi(a_{i}|s_{t},\theta_{a}).$ (11)
-Both are encoded and processed as described above when discussing ToMTrack. Note that we have ignored any hidden or memory state in the above formulation for notational convenience.
-Results in Table 8 show that ToMTrack-SAD performs worse than CNN-TrXL or other baselines.
-We speculate that including greedy actions as input is uninformative in an environment with many actions like the YLE.
+We show an example game an agent wins in self-play in Figure 13 and one where it loses in Figure 14.
+Appendix H 2 Player Cross-Play Matrices
+Figure 15: Cross-play scores for 2 player self-play as featured in Table 2 ($\alpha=0.05$).
+Figure 16: Cross-play scores for 2 player HE as featured in Table 2 ($\alpha=0.07$, $\lambda_{\text{GAE}}=0.85$).
+Figure 17: Cross-play scores for 2 player OP as featured in Table 2.
+Figure 18: Cross-play scores for 2 player OP ($\Phi_{\text{rot}}$ only) as featured in Table 2.
+Figure 19: Cross-play scores for 2 player OP ($\Phi_{\text{colour}}$ only) as featured in Table 2.
+Figure 20: Cross-play scores for 2 player OBL level 1 as featured in Table 2.
+Figure 21: Cross-play scores for 2 player OBL level 2 as featured in Table 2.
+Figure 22: Cross-play scores for 2 player OBL level 3 as featured in Table 2.
+Figure 23: Cross-play scores for 2 player OBL level 4 as featured in Table 2.
+Figure 24: Cross-play scores for 2 player OBL level 5 as featured in Table 2.
+We show cross-play matrices for self-play (Figure 15, HE (Figure 16), Other-Play (for $\Phi$ in Figure 17, for $\Phi_{\text{rot}}$ in Figure 18 and for $\Phi_{\text{colour}}$ in Figure 19) and OBL (level 1 Figure 20, level 2 Figure 21, level 3 Figure 22, level 4 Figure 23 and level 5 Figure 24).
+Appendix I Training Curves
+Figure 25: SP training return for all two player OP results as shown in Table 3.
+Figure 26: SP training curves for for two player OBL as shown in Table 2.
+Figure 27: Training curves for all belief model levels for two player OBL as shown in Table 2.
+We show relevant training curves in Figure 25, Figure 26 and Figure 27.
+All curves show convergence.
+As discussed in the main text, OP with $\alpha=0.08$ never learns any meaningful cooperation abilities and therefore shows a flat learning curve.
+More insightful are the OBL learning curves.
+We can see that the belief model improves hugely from level 0 to level 1 which is expected given that the belief model now is trained on competent self-play behaviour.
+After this however, we surprisingly find that future belief models all level out at a final cross-entropy loss of about $0.36$.
+Consequently, higher levels of OBL policies also start to saturate at similar final return levels in self-play.